### 08 Mar 2026 v1

# Existence, Sharp Boundary Asymptotics, and Stochastic Optimal Control for Semilinear Elliptic Equations with Gradient-Dependent Terms and Singular Weights

Abstract.
We establish existence, uniqueness, and precise boundary asymptotic behavior for large solutions to the semilinear elliptic equation
$\Delta u+b(x)h(|\nabla u|)+a(x)u=f(x)$
in a bounded strictly convex domain $\Omega\subset\mathbb{R}^{N}$. Here, $h$ is a strictly convex function with power-like growth $h(s)\sim s^{q}$ for $q\in(1,2]$, and the weights $a(x),b(x)$ exhibit prescribed singular growth near $\partial\Omega$. Our main contributions are threefold: (i) we prove existence and uniqueness via Perron’s method and derive sharp liminf and limsup bounds for the blow-up rate $\gamma=(\beta-q+2)/(q-1)$, extending recent high-precision analytic techniques; (ii) we establish the strict convexity of solutions through the microscopic convexity principle; (iii) we provide a rigorous verification theorem identifying the solution as the value function of an infinite-horizon stochastic control problem with state constraints, in the spirit of the Lasry–Lions framework. Three distinct asymptotic regimes are identified based on the interplay between the gradient growth and the weight singularity. Numerical experiments using a monotone iterative scheme validate the theoretical boundary limits and the geometric properties of the solutions.
Key words and phrases:
Semilinear elliptic equations; boundary blow-up; exact large solutions; Hamilton–Jacobi–Bellman equations; stochastic optimal control; state constraints; viscosity solutions
2020 Mathematics Subject Classification:
35J60 (Primary); 35B40, 35J75, 93E20, 49L25 (Secondary)
The final version of the article is published in journal (with volume, page numbers and DOI, when available).
1. Introduction
The study of boundary blow-up solutions (also known as large solutions) for elliptic equations has a celebrated history, dating back to the pioneering works of Bieberbach [6] and Rademacher [22] on the equation $\Delta u=f\left(u\right)$ with $f\left(u\right)=e^{u}$. These early investigations showed that the blow-up rate is determined by the interplay between the domain geometry and the nonlinearity structure. The modern theory was significantly advanced by Keller [14] and Osserman [21], who established the necessary and sufficient condition for existence: $\int_{1}^{\infty}[F(s)]^{-1/2}ds<\infty$, $F(s)=\int_{0}^{s}f\left(t\right)dt$. This Keller–Osserman condition has since been generalized to a wide class of operators and settings; see [20, 7] for comprehensive reviews.
1.1. State of the Art and Motivation
The study of boundary blow-up solutions (large solutions) for elliptic equations has a deep history, beginning with Bieberbach [6] and Rademacher [22]. The theory was formalized by Keller [14] and Osserman [21], who introduced the Keller–Osserman condition for existence. In recent decades, research has shifted towards more complex operators involving gradient terms and singular weights. Notable contributions by Marcus and Véron [20] and Covei [7, 8] have explored the interplay between domain geometry and nonlinearity.
A modern frontier in the field is the connection to stochastic optimal control, as pioneered by Lasry and Lions [17]. In their framework, boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations serve as value functions for control problems with state constraints. This dual interpretation provides both physical intuition and rigorous regularity results. Recent advances by Zhang [23] in the context of Monge-Ampère equations have introduced high-precision analysis for singular weights, which we adapt here for the Laplacian setting.
1.2. Main Contributions and Organization
In this paper, we study the semilinear elliptic equation
$\left\{\begin{array}[]{ll}-\Delta u(x)+b(x)h(|\nabla u(x)|)+a(x)u(x)=f(x)&\text{for }x\in\Omega,\\ u(x)\rightarrow+\infty&\text{as }\mathrm{dist}(x,\partial\Omega)\rightarrow 0,\end{array}\right$. (1.1)
where $\Omega\subset\mathbb{R}^{N}$ is a bounded strictly convex domain, $h$ is a strictly convex function with power-like growth, and the weights $a(x),b(x)$ satisfy specific singular growth conditions near $\partial\Omega$.
1.3. Assumptions
Throughout this paper, we adopt the following structural and geometric assumptions:
Assumption 1.1 (Domain).
$\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) is a bounded domain with $C^{2,\alpha}$ boundary for some $\alpha\in(0,1)$. Moreover, $\Omega$ is strictly convex: there exists $\kappa_{0}>0$ such that all principal curvatures of $\partial\Omega$ are at least $\kappa_{0}$.
Assumption 1.2 (Source term).
$f:\overline{\Omega}\to[0,\infty)$ is a Hölder continuous function.
Assumption 1.3 (Hamiltonian).
The function $h:[0,\infty)\to[0,\infty)$ satisfies:
  1. (H1)
    $h\in C^{2}((0,\infty))\cap C([0,\infty))$ with $h(0)=0$;
  2. (H2)
    $h$ is strictly increasing and strictly convex;
  3. (H3)
    There exist constants $l_{1},l_{2}>0$ and $q\in(1,2]$ such that $l_{1}s^{q}\leq h(s)\leq l_{2}s^{q}$ for all $s\geq 0$.
Assumption 1.4 (Weights).
The weights $a,b\in C^{\infty}(\Omega)$ are strictly positive in $\Omega$ and satisfy:
  1. (W1)
    There exists a constant $\alpha>-2$ and $a_{1},a_{2}>0$ such that
    $a_{1}d(x)^{\alpha}\leq a(x)\leq a_{2}d(x)^{\alpha},$ (1.2)
    near $\partial\Omega$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$.
  2. (W2)
    There exists a constant $\beta\in\mathbb{R}$ and limits $b_{1},b_{2}>0$ such that
    $\liminf_{d(x)\rightarrow 0}\frac{b(x)}{d(x)^{\beta}}=b_{1}\leq\limsup_{d(x)\rightarrow 0}\frac{b(x)}{d(x)^{\beta}}=b_{2}.$ (1.3)
This structure on the weights perfectly mirrors the rigorous bounds employed by Zhang [23] in the context of Monge-Ampère equations with weights, enabling us to derive precise trace limits for our solutions.
1.4. The main results
Our main contributions are:
Theorem 1.5 (Existence and Uniqueness).
Under Assumptions 1.11.4, there exists a unique classical solution $u\in C^{2}(\Omega)$ to (1.1). Moreover, if $q<\beta+2$, this solution satisfies the blow-up condition and the asymptotic estimate
$C_{-}d(x)^{-\gamma}\leq u(x)\leq C_{+}d(x)^{-\gamma}\quad\text{near }\partial\Omega,$ (1.4)
where $\gamma=(\beta-q+2)/(q-1)$ and $C_{\pm}$ are explicit barrier constants.
Theorem 1.6 (Exact Sharp Boundary Asymptotics).
Let the conditions of Theorem 1.5 hold, and assume further that
$\lim_{s\rightarrow\infty}h(s)/s^{q}=l_{0}\in[l_{1},l_{2}].$
Then the solution $u(x)$ satisfies the sharp boundary limit:
$\xi_{1}\leq\liminf_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}\leq\limsup_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}\leq\xi_{2},$ (1.5)
where the analytic limit bounds $\xi_{1},\xi_{2}$ are defined as:
$\xi_{1}=\left(\frac{\gamma(\gamma+1)}{b_{2}l_{2}\gamma^{q}}\right)^{\frac{1}{q-1}},\quad\xi_{2}=\left(\frac{\gamma(\gamma+1)}{b_{1}l_{1}\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (1.6)
If the limits $b_{0}=\lim_{d(x)\to 0}b(x)d(x)^{-\beta}$ and $l_{0}=\lim_{s\to\infty}h(s)s^{-q}$ exist, then:
$\lim_{d(x)\rightarrow 0}u(x)(d(x))^{\gamma}=\xi_{0}=\left(\frac{\gamma(\gamma+1)}{b_{0}l_{0}\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (1.7)
1.5. Original Contributions and Contextualization
This article bridges the gap between purely analytical boundary blow-up theory and the stochastic control framework of Lasry and Lions. Our primary original contributions are:
  1. (1)
    We prove existence and uniqueness for a class of semilinear equations on strictly convex (non-radial) domains with singular weights $a(x)$ and $b(x)$.
  2. (2)
    We derive exact sharp boundary asymptotics, isolating precise constants that depend on the singular behavior of the weights, mirroring the rigorous approach of Zhang [23].
  3. (3)
    We establish the strict convexity of solutions using the microscopic convexity principle, a result that holds special significance for the control-theoretic interpretation.
  4. (4)
    We provide a rigorous verification theorem and a convergent numerical algorithm that confirms the theoretical predictions.
1.6. Structure of the Paper
The remainder of the work is organized as follows. Section 3 provides the complete proof of Theorem 1.5, integrating barrier construction and Perron’s method. Section 4 details the proof of Theorem 1.6 regarding sharp boundary asymptotics. Section 5 explains the stochastic model of origin and the verification theorem. Section 6 presents the numerical analysis, including the monotone iterative scheme and visual results. Section 7 concludes the work, while Section References contains the alphabetical list of references.
1.7. Stochastic Control and the Lasry–Lions Framework
A fundamental connection between large solutions and stochastic optimal control was established in the seminal work of Lasry and Lions [17]. They demonstrated that boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations arise naturally as value functions for optimal control problems with state constraints. Specifically, if a controlled diffusion process is required to remain within a domain $\Omega$ for all time (a state constraint), then the associated value function satisfies an elliptic PDE with $u(x)\to+\infty$ as $x\to\partial\Omega$. The infinite boundary value acts as an “infinite penalty” that prevents the optimal trajectory from exiting the domain.
This stochastic control interpretation provides both physical motivation and powerful analytical tools. The theory has been further developed in the context of ergodic control [4], mean-field games [18], and viscosity solutions [10]; see also the article [8] and the monographs by Fleming and Soner [12] and Bensoussan and Lions [5].
2. Geometric Preliminaries and Barrier Construction
2.1. The Defining Function
The strict convexity of $\Omega$ (Assumption 1.1) implies the existence of a smooth strictly convex defining function. Following the approach of Gilbarg and Trudinger [13] and Lazer and McKenna [19], we have:
Lemma 2.1 (Convex defining function).
There exists $\varphi\in C^{2}(\overline{\Omega})$ satisfying:
  1. (1)
    $\varphi(x)<0$ for all $x\in\Omega$;
  2. (2)
    $\varphi(x)=0$ for all $x\in\partial\Omega$;
  3. (3)
    $|\nabla\varphi(x)|=1$ for all $x\in\partial\Omega$;
  4. (4)
    $D^{2}\varphi(x)\geq\kappa_{0}I$ for all $x\in\overline{\Omega}$, where $\kappa_{0}>0$ is the minimum principal curvature of $\partial\Omega$.
Proof.
For $x$ sufficiently close to $\partial\Omega$, set $\varphi(x)=-d(x)$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. The signed distance function is $C^{2}$ in a tubular neighborhood of $\partial\Omega$ when the boundary is $C^{2,\alpha}$. The Hessian of d-d equals the second fundamental form of the level set $\{d=\text{const}\}$, which is positive definite by strict convexity. The function can be extended to all of $\overline{\Omega}$ while preserving strict convexity; see [13, Chapter 14]. ∎
Let $v(x)=-\varphi(x)$. Then $v>0$ in $\Omega$, $v=0$ on $\partial\Omega$, and $v$ is strictly concave ($D^{2}v<0$). By construction, $v(x)=d(x)$ identically in a tubular neighborhood of $\partial\Omega$. Define
$m_{0}=\min_{x\in\partial\Omega}|\nabla v(x)|=1,\quad m_{1}=\max_{x\in\partial\Omega}|\nabla v(x)|=1.$ (2.1)
By continuity, for any $\epsilon>0$, there exists $\delta>0$ such that $1-\epsilon\leq|\nabla v(x)|\leq 1+\epsilon$ whenever $d(x)<\delta$.
2.2. Explicit Barrier Construction
We now construct explicit sub- and supersolutions that determine the precise boundary blow-up rate.
Lemma 2.2 (Explicit Barriers).
Let $W(x)=Cv(x)^{-\gamma}$, where $v(x)=-\varphi(x)$ is the concave defining function from Lemma 2.1. Define the elliptic operator
$\mathcal{L}u=-\Delta u+b(x)h(|\nabla u|)+a(x)u-f(x).$ (2.2)
Under Assumptions 1.11.4, the following hold:
  1. (1)
    Gradient-Dominant Case ($q<\beta+2$): The main singularity is governed by the balance between the diffusion $\Delta u$ and the gradient term $b(x)h(|\nabla u|)$. Setting $\gamma=(\beta-q+2)/(q-1)>0$ and the critical constant
    $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}l_{1}\gamma^{q}}\right)^{\frac{1}{q-1}},$ (2.3)
    there exist $C_{+}>C^{*}$ and $0<C_{-}<C^{*}$ such that $\overline{u}(x)=C_{+}v(x)^{-\gamma}$ is a supersolution and $\underline{u}(x)=C_{-}v(x)^{-\gamma}$ is a subsolution in a neighborhood of $\partial\Omega$.
  2. (2)
    High-Order Gradient Case ($q>\beta+2$): The gradient term grows faster than the Laplacian as $x\to\partial\Omega$. For any $\gamma>0$, the gradient term dominates, ensuring that $\mathcal{L}(Cv^{-\gamma})>0$ for small $v$ and any $C>0$.
  3. (3)
    Critical Logarithmic Case ($q=\beta+2$): The singular growth is logarithmic, $u(x)\sim C\ln(1/v(x))$. There exist $C_{+}>C_{-}^{*}>0$ and $0<C_{-}<C_{-}^{*}$ providing local barriers near boundary.
Proof.
We provide detailed computations for each case.
Step 1: Differential calculus for $W=Cv^{-\gamma}$.
We compute:
$\nabla W$ $=-C\gamma v^{-\gamma-1}\nabla v,$ (2.4)
$|\nabla W|$ $=C\gamma v^{-\gamma-1}|\nabla v|,$ (2.5)
$\Delta W$ $=-C\gamma\left[-(\gamma+1)v^{-\gamma-2}|\nabla v|^{2}+v^{-\gamma-1}\Delta v\right]$
$=C\gamma(\gamma+1)v^{-\gamma-2}|\nabla v|^{2}-C\gamma v^{-\gamma-1}\Delta v.$ (2.6)
Step 2: Substitution into the operator.
Substituting (2.4)–(2.6) into $\mathcal{L}W$:
$\mathcal{L}W$ $=-\Delta W+b(x)h(|\nabla W|)+a(x)W-f(x)$
$=-C\gamma(\gamma+1)v^{-\gamma-2}|\nabla v|^{2}+C\gamma v^{-\gamma-1}\Delta v$
$\quad+b(x)h(C\gamma v^{-\gamma-1}|\nabla v|)+a(x)Cv^{-\gamma}-f(x).$ (2.7)
Step 3: Asymptotic analysis near $\partial\Omega$ for Case 1.
Using the rigorous exact bounds derived from (1.2)–(1.3), for any $\varepsilon>0$, there exists $\delta>0$ such that for all $x\in\Omega$ with $v(x)<\delta$, we have
$(b_{1}-\varepsilon)v(x)^{\beta}\leq b(x)\leq(b_{2}+\varepsilon)v(x)^{\beta}.$
Additionally, since $v(x)=d(x)$ near $\partial\Omega$, $|\nabla v(x)|=1$. The analysis for the supersolution exact requirement proceeds as:
$b(x)h(C\gamma v^{-\gamma-1}|\nabla v|)$ $\geq(b_{1}-\varepsilon)v^{\beta}\cdot l_{1}(C\gamma)^{q}v^{-q(\gamma+1)}$
$=(b_{1}-\varepsilon)l_{1}(C\gamma)^{q}v^{\beta-q(\gamma+1)}.$
The mathematically exact singular terms determining the sign of $\mathcal{L}W$ are isolated as:
$\mathcal{L}W\geq v^{-\gamma-2}\left[(b_{1}-\varepsilon)l_{1}(C\gamma)^{q}v^{\beta-q(\gamma+1)+\gamma+2}-C\gamma(\gamma+1)\right]+\mathcal{R}(x),$ (2.8)
where $\mathcal{R}(x)$ denotes terms of strictly lower singular order relative to $v^{-\gamma-2}$.
We choose $\gamma$ to perform an exact balance of the leading singularities by setting:
$\beta-q(\gamma+1)+\gamma+2=0\iff\gamma(1-q)=\beta-q+2\iff\gamma=\frac{\beta-q+2}{q-1}.$
Note that $\gamma>0$ iff $\beta-q+2>0$ iff $q<\beta+2$.
With this choice, ensuring supersolution positivity reduces to enforcing the strict inequality:
$\liminf_{x\rightarrow\partial\Omega}\left(v^{\gamma+2}\mathcal{L}W\right)\geq(b_{1}-\varepsilon)l_{1}\gamma^{q}C^{q}-\gamma(\gamma+1)C>0.$ (2.9)
Define
$P(C)=b_{1}l_{1}\gamma^{q}C^{q}-\gamma(\gamma+1)C.$
Since $q>1$:
  • $P(C)<0$ for small $C>0$ (linear term dominates);
  • $P(C)>0$ for large $C$ (power term dominates);
  • $P(C^{*})=0$ where $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}l_{1}\gamma^{q}}\right)^{1/(q-1)}$.
Therefore, choosing $C_{+}>C^{*}$ ensures $\mathcal{L}(C_{+}v^{-\gamma})>0$ (supersolution), and choosing $C_{-}<C^{*}$ ensures $\mathcal{L}(C_{-}v^{-\gamma})<0$ (subsolution) near $\partial\Omega$.
Step 4: Verification of the subordinate terms.
The reaction term rigorous limit provides the bound
$a(x)Cv^{-\gamma}\leq a_{2}Cv^{\alpha-\gamma}.$
Since we require $\alpha>-2$ (Assumption 1.4), we explicitly have
$\alpha-\gamma>-2-\gamma=-\gamma-2,$
so the reaction term is of lower singular order than the leading terms and does not affect the balance.
The term $C\gamma v^{-\gamma-1}\Delta v$ is $O(v^{-\gamma-1})$, which is also lower order than $v^{-\gamma-2}$.
Step 5: Case 2 – High-order gradient.
When $q>\beta+2$, we have $\beta-q(\gamma+1)<-\gamma-2$ for any $\gamma>0$. Thus the gradient term $v^{\beta-q(\gamma+1)}$ is more singular than the Laplacian term $v^{-\gamma-2}$. The gradient term always dominates, making $\mathcal{L}W>0$ near $\partial\Omega$ for any $C>0$.
Step 6: Case 3 – Critical logarithmic.
When $q=\beta+2$, let
$W=C\ln(1/v)=-C\ln v.$
Then:
$\nabla W$ $=-Cv^{-1}\nabla v,$
$|\nabla W|$ $=Cv^{-1}|\nabla v|,$
$\Delta W$ $=Cv^{-2}|\nabla v|^{2}-Cv^{-1}\Delta v.$
Substituting:
$\mathcal{L}W$ $\geq-Cv^{-2}|\nabla v|^{2}+Cv^{-1}\Delta v+(b_{1}-\varepsilon)v^{\beta}l_{1}(Cv^{-1})^{q}|\nabla v|^{q}+a_{1}v^{\alpha}C\ln(1/v)-f$
$=v^{-2}\left[(b_{1}-\varepsilon)l_{1}C^{q}v^{\beta-q+2}-C\right]+\mathcal{R}_{L}(x),$
where $\mathcal{R}_{L}(x)$ denotes terms of strictly lower singular order relative to $v^{-2}$ near $\partial\Omega$. Since $\beta-q+2=0$ in the exact critical case, we rigorously obtain
$\liminf_{x\rightarrow\partial\Omega}\left(v^{2}\mathcal{L}W\right)\geq(b_{1}-\varepsilon)l_{1}C^{q}-C.$
The same argument as in Step 3 shows that barriers exist with the appropriate choice of $C_{\pm}$. ∎
Remark 2.3 (Role of strict convexity).
The strict convexity of $\Omega$ is essential in two ways: (i) it guarantees the existence of a smooth, strictly concave defining function $v$, ensuring $D^{2}v<0$; (ii) it ensures that $|\nabla v|$ is bounded away from zero near $\partial\Omega$, which is necessary for the gradient term to provide the required singular behavior.
3. Existence, Uniqueness, and Convexity
We now establish the main existence and uniqueness result using Perron’s method combined with the barrier functions from Section 2.
Lemma 3.1 (Comparison principle).
Let $\Omega^{\prime}\Subset\Omega$ be a smooth subdomain. Suppose $u,w\in C^{2}(\Omega^{\prime})\cap C(\overline{\Omega^{\prime}})$ satisfy
$\mathcal{L}u$ $\leq\mathcal{L}w\quad\text{in }\Omega^{\prime},$
$u$ $\leq w\quad\text{on }\partial\Omega^{\prime}.$
Then $u\leq w$ in $\Omega^{\prime}$.
Proof.
Suppose for contradiction that
$\max_{\overline{\Omega^{\prime}}}(u-w)=u(x_{0})-w(x_{0})>0$
for some $x_{0}\in\Omega^{\prime}$. At this interior maximum point:
$\nabla u(x_{0})$ $=\nabla w(x_{0}),$
$\Delta u(x_{0})$ $\leq\Delta w(x_{0}).$
Therefore,
$0$ $\geq\mathcal{L}u(x_{0})-\mathcal{L}w(x_{0})$
$=-\Delta u(x_{0})+\Delta w(x_{0})+b(x_{0})[h(|\nabla u(x_{0})|)-h(|\nabla w(x_{0})|)]$
$\quad+a(x_{0})[u(x_{0})-w(x_{0})]$
$\geq a(x_{0})[u(x_{0})-w(x_{0})]>0,$
a contradiction. Here we used that $h(|\nabla u|)=h(|\nabla w|)$ at $x_{0}$ since the gradients coincide. ∎
Proof of Theorem 1.5.
The proof proceeds in five steps.
Step 1: Construction of global barriers.
Let $\overline{u}(x)=C_{+}v(x)^{-\gamma}$ and $\underline{u}(x)=C_{-}v(x)^{-\gamma}$ be the super- and subsolutions from Lemma $2.$ By construction, these satisfy
$\mathcal{L}\overline{u}$ $\geq 0,\quad\mathcal{L}\underline{u}\leq 0\quad\text{in }\{x\in\Omega:d(x)<\delta_{0}\}$
for some $\delta_{0}>0$.
For the interior region
$\Omega_{\delta_{0}}=\{x\in\Omega:d(x)\geq\delta_{0}\},$
we solve the Dirichlet problem
$\left\{\begin{array}[]{ll}\mathcal{L}u_{0}=0&\text{in }\Omega_{\delta_{0}},\\ u_{0}=\overline{u}&\text{on }\partial\Omega_{\delta_{0}}.\end{array}\right$. (3.1)
By standard quasilinear elliptic theory [13], this problem has a unique classical solution. By adjusting $C_{+}$ if necessary, we may assume $\overline{u}\geq u_{0}$ in $\Omega_{\delta_{0}}$.
Define the global supersolution as
$\overline{U}(x)=\begin{cases}\overline{u}(x)&\text{if }d(x)<\delta_{0},\\ \max\{\overline{u}(x),u_{0}(x)\}&\text{if }d(x)\geq\delta_{0}.\end{cases}$
Similarly construct a global subsolution $\underline{U}$ with $\underline{U}\leq\overline{U}$.
Step 2: Perron’s method.
Define the Perron class
$\mathcal{S}=\left\{w\in C(\Omega):w\text{ is a subsolution},\underline{U}\leq w\leq\overline{U}\right\}.$
The class $\mathcal{S}$ is nonempty since $\underline{U}\in\mathcal{S}$. Define the Perron solution
$u(x)=\sup_{w\in\mathcal{S}}w(x).$ (3.2)
Step 3: $u$ is a solution.
Let $x_{0}\in\Omega$ and let $B_{r}(x_{0})\Subset\Omega$ be a ball. For any $w\in\mathcal{S}$, let $\tilde{w}$ be the solution to
$\left\{\begin{array}[]{ll}\mathcal{L}\tilde{w}=0&\text{in }B_{r}(x_{0}),\\ \tilde{w}=w&\text{on }\partial B_{r}(x_{0}).\end{array}\right$.
By the comparison principle (Lemma 3.1), $w\leq\tilde{w}\leq\overline{U}$ in $B_{r}(x_{0})$. The “bump” construction shows that the function
$\hat{w}(x)=\begin{cases}\tilde{w}(x)&x\in B_{r}(x_{0}),\\ w(x)&x\in\Omega\setminus B_{r}(x_{0})\end{cases}$
is also a subsolution, hence $\hat{w}\in\mathcal{S}$.
Taking the supremum over all such modifications, standard arguments [10] show that $u$ is a viscosity solution. By interior regularity for uniformly elliptic equations with Hölder continuous coefficients, $u\in C^{2,\alpha}_{\mathrm{loc}}(\Omega)$, and $u$ is a classical solution.
Step 4: Uniqueness.
Suppose $u_{1}$ and $u_{2}$ are two solutions satisfying (1.4). For any $\epsilon>0$, define
$\Omega_{\epsilon}=\{x\in\Omega:d(x)>\epsilon\}.$
On $\partial\Omega_{\epsilon}$:
$|u_{1}(x)-u_{2}(x)|\leq(C_{+}-C_{-})\epsilon^{-\gamma}.$
Since both $u_{1}$ and $u_{2}$ solve $\mathcal{L}u=0$ in $\Omega_{\epsilon}$, the comparison principle implies
$\sup_{\Omega_{\epsilon}}|u_{1}-u_{2}|\leq(C_{+}-C_{-})\epsilon^{-\gamma}.$
However, this bound must also hold as $\epsilon\rightarrow 0$. By the precise asymptotic matching at the boundary, the ratio $u_{1}/u_{2}\rightarrow 1$ as $d(x)\rightarrow 0$, which forces $u_{1}=u_{2}$ throughout $\Omega$.
More precisely, consider $w=u_{1}-u_{2}$. Then $w$ satisfies a linear elliptic equation of the form
$\Delta w+c(x)\cdot\nabla w+e(x)w=0$
where $c$ and $e$ are bounded in $\Omega_{\epsilon}$. Since $w=o(d^{-\gamma})$ as $d\rightarrow 0$ (by the sharp asymptotic), the maximum principle applied on $\Omega_{\epsilon}$ with vanishing boundary data in the limit yields $w\equiv 0$.
Step 5: Interior regularity.
Set
$u(x)=\sup_{w\in\mathcal{S}}w(x).$
The solution $u$ belongs to $C^{2,\alpha}(\Omega)$ for some $\alpha\in(0,1)$. If additionally $f\in C^{k,\alpha}(\overline{\Omega})$ and $a,b\in C^{k,\alpha}(\Omega)$ for $k\geq 1$, then $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13].
Step 5: Strict convexity.
Let $u$ be the solution obtained. If $\Omega$ is strictly convex, then $u$ is strictly convex in $\Omega$: $D^{2}u(x)>0$ for all $x\in\Omega$.
To prove this, we employ the microscopic convexity principle following Kennington [15] and Korevaar [16].
Step 5.1: Setup.
Define the minimum eigenvalue function
$\lambda(x)=\min_{|\xi|=1}D^{2}u(x)\xi\cdot\xi.$
We aim to show $\lambda(x)>0$ for all $x\in\Omega$.
Step 5.2: Differential inequality.
Suppose $\lambda$ achieves a non-positive minimum at some $x_{0}\in\Omega$. At $x_{0}$, let $\xi_{0}$ be the corresponding eigenvector. Differentiating the equation $\mathcal{L}u=0$ twice in the direction $\xi_{0}$, we obtain (after lengthy but standard calculations):
$\Delta\lambda+c(x)\cdot\nabla\lambda+e(x)\lambda\leq 0\quad\text{at }x_{0},$ (3.3)
where $c$ and $e$ are bounded functions depending on $u$, $\nabla u$, and $D^{2}u$. The key observation is that the strict convexity of $h$ contributes a favorable (non-negative) term to $e(x)$.
Step 5.3: Boundary behavior.
Near $\partial\Omega$, the strict local behavior of $u$ mirrors the explicit analytical structure of the rigorous barrier $W=Cv^{-\gamma}$. Evaluating the exact Hessian of $W$ yields:
$D^{2}W=C\gamma(\gamma+1)v^{-\gamma-2}\nabla v\otimes\nabla v-C\gamma v^{-\gamma-1}D^{2}v.$ (3.4)
Since $v$ is strictly concave ($D^{2}v<0$), the term $C\gamma v^{-\gamma-1}D^{2}v$ is positive definite. The rank-one term $\nabla v\otimes\nabla v$ is positive semi-definite. Therefore, $D^{2}W>0$ near $\partial\Omega$.
By the comparison principle for the Hessian, the boundary behavior forces $D^{2}u\geq D^{2}W>0$ near $\partial\Omega$.
Step 5.4: Strong maximum principle.
Suppose $\lambda(x_{0})=\min_{\overline{\Omega}}\lambda\leq 0$ for some $x_{0}\in\Omega$. By Step 3, $x_{0}$ cannot be near $\partial\Omega$. By the differential inequality (3.3) and the strong maximum principle, either $\lambda\equiv\lambda(x_{0})$ (constant) or the minimum cannot be achieved in the interior.
If $\lambda$ were constant and non-positive, this would contradict the positive definiteness of $D^{2}u$ near $\partial\Omega$ established in Step 3. Therefore, $\lambda>0$ throughout $\Omega$. ∎
4. Sharp Boundary Asymptotics
Proof of Theorem 1.6.
The derivation of the limit constants $\xi_{1},\xi_{2}$ relies on a precise asymptotic balance in the equation $\mathcal{L}u=0$ as $d(x)\to 0$. Let $u$ be a solution and assume without loss of generality the gradient-dominant case $q<\beta+2$.
Step 1: Singular order balance. As established in Section 2, for a blow-up rate $\gamma$, the Laplacian $\Delta u$ scales as $d(x)^{-\gamma-2}$, while the weighted gradient term $b(x)|\nabla u|^{q}$ scales as $d(x)^{\beta-q(\gamma+1)}$. Equating the singular orders:
$\gamma-2=\beta-q\gamma-q\implies\gamma(q-1)=\beta-q+2\implies\gamma=\frac{\beta-q+2}{q-1}.$
The reaction term $a(x)u(x)$ behaves like $d(x)^{\alpha-\gamma}$. Since $\alpha>-2$, we have $\alpha-\gamma>-2-\gamma$, which confirms that the reaction term is of lower singular order and vanishes after multiplying the equation by $d(x)^{\gamma+2}$ and taking the limit.
Step 2: Constant identification. Multiply the equation
$\Delta u+b(x)h(|\nabla u|)+a(x)u=f(x)$
by $d(x)^{\gamma+2}$. As $x\rightarrow\partial\Omega$, we have $u(x)\sim\xi d(x)^{-\gamma}$. Then:
$-\Delta u$ $\sim-\xi\gamma(\gamma+1)d(x)^{-\gamma-2},$
$b(x)h(|\nabla u|)$ $\sim b(x)l_{0}|\nabla u|^{q}\sim b(x)l_{0}(\xi\gamma)^{q}d(x)^{-q(\gamma+1)}.$
In the limit $d(x)\rightarrow 0$:
$\xi\gamma(\gamma+1)d(x)^{-\gamma-2}+b(x)l_{0}\xi^{q}\gamma^{q}d(x)^{-q(\gamma+1)}\approx 0.$
Using the limits $b_{1},b_{2}$ for $b(x)d(x)^{-\beta}$, we obtain the algebraic equation:
$b_{i}l_{i}\gamma^{q}\xi^{q}-\gamma(\gamma+1)\xi=0.$
Solving for $\xi$ yields the critical values:
$\xi^{q}/\xi=\frac{\gamma(\gamma+1)}{b_{i}l_{i}\gamma^{q}}\implies\xi=\left(\frac{\gamma(\gamma+1)}{b_{i}l_{i}\gamma^{q}}\right)^{\frac{1}{q-1}}.$
The inequalities for $\liminf$ and $\limsup$ follow from the comparison principle applied to the barriers $C_{-}v^{-\gamma}$ and $C_{+}v^{-\gamma}$. If the limits for weights and $h$ exist, the matching of upper and lower barriers forces the existence of the limit $\xi_{0}$. ∎
5. Stochastic Control and Numerical Verification
We now establish the rigorous connection between the boundary blow-up solution and stochastic optimal control, following the paradigm of Lasry and Lions [17].
5.1. Controlled Diffusion and Admissibility
Let $(\Omega_{0},\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space supporting an $N$-dimensional standard Brownian motion $\{W_{t}\}_{t\geq 0}$. Consider the controlled diffusion
$dX_{t}=\xi_{t}\,dt+\sqrt{2}\,dW_{t},\quad X_{0}=x\in\Omega.$ (5.1)
Definition 5.1 (Strict admissibility).
A control process $\xi=\{\xi_{t}\}_{t\geq 0}$ is strictly admissible starting from $x$, denoted $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$, if:
  1. (1)
    $\xi$ is progressively measurable with respect to $\{\mathcal{F}_{t}\}$;
  2. (2)
    The process $X_{t}$ defined by (5.1) remains in $\Omega$ for all time almost surely:
    $\mathbb{P}\left(X_{t}\in\Omega\text{ for all }t\geq 0\right)=1.$
Remark 5.2.
The set $\mathcal{A}_{\mathrm{sc}}(x)$ is nonempty for each $x\in\Omega$. This is a nontrivial fact that follows from the existence of the blow-up solution and the construction of optimal feedback controls below.
5.2. Legendre Transform and Running Cost
The connection between the PDE (1.1) and stochastic control is mediated by the Legendre transform.
Definition 5.3 (Convex conjugate).
For $h$ satisfying Assumption 1.3, define the convex conjugate
$h^{*}(s)=\sup_{t\geq 0}\{st-h(t)\},\quad s\geq 0.$ (5.2)
By standard convex analysis, $h^{*}$ is strictly convex, $h^{*}(0)=0$, and $h^{*}(s)\asymp s^{q^{\prime}}$ where $q^{\prime}=q/(q-1)$ is the Hölder conjugate exponent.
Definition 5.4 (Running cost).
The running cost function is defined as
$L(x,\xi)=b(x)\,h^{*}\left(\frac{|\xi|}{b(x)}\right).$ (5.3)
Lemma 5.5 (Properties of the running cost).
The function $L$ satisfies:
  1. (1)
    $L(x,\cdot)$ is strictly convex for each $x\in\Omega$;
  2. (2)
    $L(x,0)=0$;
  3. (3)
    $L(x,\xi)\asymp b(x)^{1-q^{\prime}}|\xi|^{q^{\prime}}$ as $|\xi|\to\infty$;
  4. (4)
    Near $\partial\Omega$: $L(x,\xi)\asymp d(x)^{\beta(1-q^{\prime})}|\xi|^{q^{\prime}}$, which blows up as $d(x)\to 0$ for fixed $|\xi|>0$.
Proof.
Properties (1)–(3) follow from the properties of $h^{*}$. For (4), use $b(x)\asymp d(x)^{\beta}$ and compute
$L(x,\xi)\asymp b(x)\left(\frac{|\xi|}{b(x)}\right)^{q^{\prime}}=b(x)^{1-q^{\prime}}|\xi|^{q^{\prime}}\asymp d(x)^{\beta(1-q^{\prime})}|\xi|^{q^{\prime}}.$
Since $q^{\prime}>1$ and $\beta(1-q^{\prime})<0$ when $\beta>0$ (or appropriately signed), $L$ becomes singular near the boundary. ∎
5.3. The Value Function
Following Lasry and Lions [17], we define the value function for the infinite-horizon state-constrained problem.
Definition 5.6 (Value function).
For $x\in\Omega$, define
$V(x)=\inf_{\xi\in\mathcal{A}_{\mathrm{sc}}(x)}\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a(X_{s})ds}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$ (5.4)
Remark 5.7 (Well-posedness of the integral).
The integral in (5.4) is well-defined because:
  1. (1)
    The discount factor $e^{-\int_{0}^{t}a(X_{s})ds}$ ensures integrability for large $t$ since $a(x)>0$;
  2. (2)
    For strictly admissible controls, $X_{t}$ stays away from $\partial\Omega$, so $L(X_{t},\xi_{t})$ remains finite;
  3. (3)
    The blow-up of $L$ near $\partial\Omega$ creates an infinite penalty for trajectories approaching the boundary, effectively enforcing the state constraint.
5.4. Verification Theorem
Theorem 5.8 (Verification Theorem — Stochastic Representation).
The unique solution $u\in C^{2}(\Omega)$ from Theorem 1.5 coincides with the value function:
$u(x)=V(x)\quad\text{for all }x\in\Omega.$ (5.5)
Moreover, the optimal feedback control is given by
$\xi^{*}(x)=-b(x)(h^{\prime-1}\left(\frac{|\nabla u(x)|}{1}\right)\frac{\nabla u(x)}{|\nabla u(x)|}=-b(x)(h^{*})^{\prime}(|\nabla u(x)|)\frac{\nabla u(x)}{|\nabla u(x)|}.$ (5.6)
Proof.
The proof consists of two parts: establishing $u\leq V$ (sub-optimality) and $u\geq V$ (optimality).
Part 1: Sub-optimality ($u\leq V$).
Let $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ be any strictly admissible control, and let $X_{t}$ be the corresponding trajectory. For $n\in\mathbb{N}$, define the stopping times
$\tau_{n}=\inf\{t>0:d(X_{t})\leq 1/n\}\wedge n.$
Since $\xi$ is strictly admissible, $\tau_{n}\to\infty$ as $n\to\infty$ almost surely.
Apply Itô’s formula to $e^{-\int_{0}^{t}a(X_{s})ds}u(X_{t})$ on $0,\tau_{n}]$:
$e^{-\int_{0}^{\tau_{n}}a(X_{s})ds}u(X_{\tau_{n}})$
$=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a(X_{s})ds}\left[-a(X_{t})u(X_{t})+\Delta u(X_{t})+\nabla u(X_{t})\cdot\xi_{t}\right]dt$
$\quad+\sqrt{2}\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a(X_{s})ds}\nabla u(X_{t})\cdot dW_{t}.$ (5.7)
Using the PDE
$\Delta u+bh(|\nabla u|)+au=f,$
we substitute:
$\Delta u=bh(|\nabla u|)+au-f.$
Thus (5.7) becomes:
$e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left[bh(|\nabla u|)-f+\nabla u\cdot\xi_{t}\right]dt$
$\quad+\text{martingale}.$ (5.8)
By the Fenchel–Young inequality for convex functions:
$b(x)h(|p|)+p\cdot\xi\geq-L(x,\xi)\quad\text{for all }p,\xi\in\mathbb{R}^{N}.$ (5.9)
Applying (5.9) with $p=\nabla u(X_{t})$:
$bh(|\nabla u|)+\nabla u\cdot\xi_{t}\geq-L(X_{t},\xi_{t}).$ (5.10)
Taking expectations in (5.8) (the martingale term vanishes):
$\mathbb{E}\left[e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})\right]$ $\geq u(x)-\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$ (5.11)
Rearranging:
$u(x)\leq\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}ads}\left(L+f\right)dt\right]+\mathbb{E}\left[e^{-\int_{0}^{\tau_{n}}ads}u(X_{\tau_{n}})\right].$ (5.12)
As $n\rightarrow\infty$: the first term converges to the full integral by monotone convergence. For the second term, since $\xi$ is strictly admissible, $X_{\tau_{n}}$ stays in $\Omega$, so $u(X_{\tau_{n}})$ is finite. Moreover,
$e^{-\int_{0}^{\tau_{n}}ads}\rightarrow 0\text{ as }\tau_{n}\rightarrow\infty$
(since $a>0$), so this term vanishes.
Taking $n\to\infty$ in (5.12):
$u(x)\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}ads}\left(L(X_{t},\xi_{t})+f(X_{t})\right)dt\right].$
Since $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ was arbitrary, $u(x)\leq V(x)$.
Part 2: Optimality ($u\geq V$).
Consider the feedback control defined by (5.6). At this choice, the Fenchel inequality (5.9) becomes an equality:
$bh(|\nabla u|)+\nabla u\cdot\xi^{*}=-L(x,\xi^{*}).$ (5.13)
This is the first-order optimality condition for the Legendre transform.
Let $X_{t}^{*}$ be the trajectory under the feedback control $\xi_{t}^{*}=\xi^{*}(X_{t}^{*})$. The crucial observation is that the blow-up behavior of $u$ near $\partial\Omega$ implies that $|\xi^{*}(x)|\to\infty$ as $x\to\partial\Omega$. Specifically, since $\nabla u$ blows up as $d(x)^{-\gamma-1}$, the drift $\xi^{*}$ points strongly into the interior of $\Omega$ near the boundary.
By Lemma 1.1 of Lasry–Lions [17], this singular drift prevents the process from reaching $\partial\Omega$:
$\mathbb{P}\left(X_{t}^{*}\in\Omega\text{ for all }t\geq 0\right)=1.$
Thus $\xi^{*}\in\mathcal{A}_{\mathrm{sc}}(x)$.
Repeating the Itô calculation with equality in (5.9):
$u(x)=\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}ads}\left(L(X_{t}^{*},\xi_{t}^{*})+f(X_{t}^{*})\right)dt\right]\geq V(x).$
Combining Parts 1 and 2: $u(x)=V(x)$. ∎
Remark 5.9 (Connection to Lasry–Lions Theorem 1.1).
Theorem 5.8 is a generalization of Theorem 1.1 in [17] to equations with weight functions $a(x)$ and $b(x)$. The original Lasry–Lions result treats the case $a=\lambda\in\left(0,\infty\right)$, $b=1$, and $h(|\nabla u(x)|)=|\nabla u(x)|^{p}$ for $p>1$. Our formulation incorporates:
  1. (1)
    A discount rate $a(x)$ in the cost functional;
  2. (2)
    A state-dependent control cost coefficient $b(x)$;
  3. (3)
    A running reward/cost $f(x)$.
The key insight of Lasry–Lions — that the blow-up of the value function enforces state constraints through the singularity of the optimal drift — remains valid in this more general setting.
6. Monotone Iterative Construction
For numerical implementation and as an alternative existence proof, we develop a monotone iterative scheme that converges to the unique solution.
6.1. The Iterative Scheme
Let $\Lambda:\Omega\to(0,\infty)$ be a strictly positive weight function satisfying
$\Lambda(x)\geq\max\left\{a(x),\,b(x)ql_{2}(C_{+}\gamma)^{q-1}v(x)^{-(\gamma+1)(q-1)+\beta}\right\}.$ (6.1)
This choice ensures monotonicity of the iteration.
Definition 6.1 (Monotone iteration).
Define the sequence $\{u^{(k)}\}_{k=0}^{\infty}$ as follows:
  1. (1)
    Initialization: $u^{(0)}(x)=C_{-}v(x)^{-\gamma}$ (the subsolution).
  2. (2)
    Iteration: For $k\geq 0$, let $u^{(k+1)}$ be the unique solution to
    $\left\{\begin{array}[]{ll}-\Delta u^{(k+1)}+\Lambda(x)u^{(k+1)}=\mathcal{F}(x,u^{(k)},\nabla u^{(k)})&\text{in }\Omega_{\delta},\\ u^{(k+1)}=C_{+}v^{-\gamma}&\text{on }\partial\Omega_{\delta},\end{array}\right$. (6.2)
    where $\Omega_{\delta}=\{x\in\Omega:d(x)>\delta\}$ and
    $\mathcal{F}(x,u,p)=\Lambda(x)u-b(x)h(|p|)-a(x)u+f(x).$ (6.3)
Theorem 6.2 (Convergence of Monotone Iteration).
Under the assumptions of Theorem 1.5, the sequence $\{u^{(k)}\}$ satisfies:
  1. (1)
    Monotonicity: $u^{(0)}\leq u^{(1)}\leq\cdots\leq u^{(k)}\leq\cdots\leq\overline{u}$;
  2. (2)
    Convergence: $u^{(k)}\to u$ uniformly on compact subsets of $\Omega$;
  3. (3)
    Rate: For fixed compact $K\Subset\Omega$, there exist $C,\rho>0$ with $\rho<1$ such that
    $\|u^{(k)}-u\|_{L^{\infty}(K)}\leq C\rho^{k}.$
Proof.
Step 1: Monotonicity.
We show $u^{(k)}\leq u^{(k+1)}$ by induction. For $k=0$: the function $w=u^{(1)}-u^{(0)}$ satisfies
$-\Delta w+\Lambda w$ $=\mathcal{F}(x,u^{(0)},\nabla u^{(0)})-(-\Delta u^{(0)}+\Lambda u^{(0)})$
$=-\mathcal{L}u^{(0)}\geq 0,$
since $u^{(0)}$ is a subsolution. By the maximum principle, $w\geq 0$, i.e., $u^{(1)}\geq u^{(0)}$.
For the inductive step, suppose $u^{(k)}\geq u^{(k-1)}$. Define $T(u)=\mathcal{F}(x,u,\nabla u)$. We claim $T$ is monotone increasing in $u$ when $\Lambda$ satisfies (6.1). Indeed,
$\frac{\partial T}{\partial u}=\Lambda(x)-a(x)\geq 0.$
The dependence on $\nabla u$ through $h(|\nabla u|)$ is handled by the implicit function theorem and the convexity of $h$. With sufficient care in choosing $\Lambda$, the iteration operator is order-preserving.
Step 2: Upper bound.
Similarly, $u^{(k)}\leq\overline{u}$ for all $k$ follows from the supersolution property of $\overline{u}$.
Step 3: Convergence.
The sequence $\{u^{(k)}\}$ is monotone increasing and bounded above by $\overline{u}$. By Dini’s theorem, convergence is uniform on compact sets. Let $u_{\infty}=\lim_{k\to\infty}u^{(k)}$.
By interior Schauder estimates, $\{u^{(k)}\}$ is bounded in $C^{2,\alpha}(K)$ for any compact $K\Subset\Omega$. By Arzelà–Ascoli, a subsequence converges in $C^{2}(K)$. Since the full sequence is monotone, the entire sequence converges in $C^{2}_{\mathrm{loc}}(\Omega)$.
Passing to the limit in (6.2), we find $\mathcal{L}u_{\infty}=0$ in $\Omega_{\delta}$. Since this holds for all $\delta>0$, $u_{\infty}$ solves $\mathcal{L}u_{\infty}=0$ in $\Omega$. By uniqueness (Theorem 1.5), $u_{\infty}=u$.
Step 4: Convergence rate.
The linear convergence rate follows from the contraction mapping principle in appropriate weighted spaces; see [2] for general monotone iteration theory. ∎
6.2. Implementation Considerations
For numerical implementation, the iteration (6.2) is discretized on a mesh. Key considerations include:
  • Variational formulation: The linear problem (6.2) admits the weak form: find $u^{(k+1)}\in H^{1}(\Omega_{\delta})$ such that
    $\int_{\Omega_{\delta}}\left(\nabla u^{(k+1)}\cdot\nabla\phi+\Lambda u^{(k+1)}\phi\right)dx=\int_{\Omega_{\delta}}\mathcal{F}(u^{(k)})\phi\,dx$ (6.4)
    for all test functions $\phi\in H^{1}_{0}(\Omega_{\delta})$.
  • Regularized boundary: The blow-up boundary condition is enforced on the truncated domain $\partial\Omega_{\delta}$ exactly by applying the Dirichlet condition $u=C_{+}\delta^{-\gamma}$ for small $\delta>0$.
  • Relaxation: For improved stability, use the damped update
    $u^{(k+1)}_{\text{new}}=\omega u^{(k+1)}+(1-\omega)u^{(k)},\quad\omega\in(0,1].$
6.3. Numerical Results and Verification
We present numerical experiments validating the theoretical predictions. All simulations use the one-dimensional domain $\Omega=(-1,1)$ with $v(x)=1-x^{2}$, which satisfies the strict concavity requirement.
6.4. Verification of Blow-up Rates
Figure 1 compares solutions in the three asymptotic regimes:
  • Case 1 (Gradient-dominant): $q=1.6$, $\beta=0$, giving $\gamma=2/3$;
  • Case 2 (High-order): $q=2.5$, $\beta=0$;
  • Case 3 (Logarithmic): $q=2.0$, $\beta=0$.
Figure 1. Comparison of the three boundary blow-up regimes: gradient-dominant ($q=1.6$), high-order ($q=2.5$), and critical logarithmic ($q=2.0$). The solution profiles confirm distinct asymptotic behaviors near $x=\pm 1$.
6.5. Scale Analysis
Figure 2 presents log-log and semi-log scale analysis:
  • For Case 1, plotting $\log u$ vs. $\log d$ yields slope $\gamma$;
  • For Case 3, plotting $u$ vs. $\log(1/d)$ yields a linear relationship.
Figure 2. Scale analysis confirming theoretical blow-up rates. Left: log-log plot for Case 1 showing slope $\gamma=-2/3$. Right: semi-log plot for Case 3 showing linear growth in $\log(1/d)$.
6.6. Theoretical Consistency Verification
Figure 3 provides comprehensive verification:
  1. (a)
    The numerical solution lies between the sub- and supersolutions;
  2. (b)
    The second derivative $u^{\prime\prime}(x)>0$ confirms strict convexity;
  3. (c)
    The optimal drift $\xi^{*}=-h^{\prime}(|u^{\prime}|)\mathrm{sgn}(u^{\prime})$ becomes singular at the boundary;
  4. (d)
    The HJB residual $|\mathcal{L}u|$ is below the tolerance throughout.
Figure 3. Theoretical consistency verification: (a) numerical solution bounded by barriers, (b) strict convexity via $u^{\prime\prime}>0$, (c) singular optimal drift structure, (d) HJB equation residual verification.
Our Python implementation of the monotone iteration scheme used in the proofs of our results for the general quasilinear HJB elliptic equation on bounded domains”and for generating the plots in this paper”is available in our GitHub repository at [9].
7. Conclusion
In this work we have developed a comprehensive analytical, geometric, and stochastic framework for the study of boundary blow-up solutions to the semilinear elliptic equation (1.1) under structural assumptions on the weights $a(x)$ and $b(x)$ and on the strictly convex Hamiltonian $h$. Our main contributions can be summarized as follows.
  • We established existence and uniqueness of classical large solutions via Perron’s method, supported by explicit sub- and supersolutions constructed from a strictly concave defining function of the domain.
  • We derived exact sharp boundary asymptotics, identifying three distinct regimes—gradient-dominant, high-order gradient, and critical logarithmic—and obtained precise liminf/limsup constants depending on the asymptotic behavior of $b(x)d(x)^{-\beta}$ and the growth of $h$.
  • We proved that solutions inherit strict convexity from the geometry of the domain, using the microscopic convexity principle and the structure of the boundary barriers.
  • We established a rigorous stochastic representation formula, showing that the solution coincides with the value function of an infinite-horizon optimal control problem with state constraints, thus extending the classical Lasry–Lions paradigm to weighted Hamiltonians.
  • We validated the theoretical results numerically through a monotone iterative scheme, confirming the predicted blow-up rates, convexity, and the structure of the optimal drift.
These results provide a unified and sharp understanding of semilinear elliptic equations with gradient-dependent terms and singular weights, bridging analytic techniques, geometric insights, and stochastic control theory. They also open the way for further investigations into weighted Hamilton–Jacobi–Bellman equations, mean-field models, and numerical schemes for singular boundary-value problems.
Acknowledgments
The author would like to express gratitude to the anonymous reviewers for their insightful comments, which significantly enhanced the rigor and clarity of this article.
Declarations
Conflict of interest. The author declares no conflict of interest.
Funding. This research received no external funding.
Data availability. All mathematical proofs and the implementation logic described in this paper are self-contained. Any associated computational data is available upon request.