### 22 Mar 2026 v2

# Existence, Sharp Boundary Asymptotics, and Stochastic Optimal Control for Semilinear Elliptic Equations with Gradient-Dependent Terms and Singular Weights: Theory, Economic Applications, and Numerical Simulations

Abstract.
This article develops a rigorous general framework for a class of semilinear elliptic equations with gradient-dependent nonlinearities and singular weights posed in strictly convex domains. We study large solutions to
$\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u=f(x)\quad\text{in }\Omega\subset\mathbb{R}^{N},$
where $h$ is strictly convex with power-type growth $h(s)\sim s^{q}$ for $q\in(1,2]$, and the coefficients $a(x)$ and $b(x)$ exhibit prescribed singular behavior near $\partial\Omega$. The main contributions are fourfold. First, we establish existence and uniqueness of large solutions via Perron’s method and derive sharp boundary asymptotics, including optimal liminf–limsup estimates for the blow-up rate $\gamma=(\beta-q+2)/(q-1)$, thereby extending recent high-precision analytic techniques. Second, we prove the strict convexity of solutions by means of the microscopic convexity principle, revealing new geometric properties induced by the interplay between gradient growth and weight singularity. Third, we provide a verification theorem showing that the solution coincides with the value function of an infinite-horizon stochastic optimal control problem with non-violable state constraints, in the spirit of the Lasry–Lions approach. Fourth, we develop a comprehensive section on applications in Operational Research and Management Science, including real-world case studies in inventory management, portfolio risk control, and supply chain optimization, and we explore interdisciplinary connections with physics through boundary layer theory and diffusion processes. Numerical experiments for both the one-dimensional ($N=1$) and two-dimensional ($N=2$) cases, based on a monotone iterative scheme, confirm the theoretical boundary behavior and validate the predicted geometric structure. All Python implementation codes are provided in the Appendix.
Key words and phrases:
Semilinear elliptic equations; boundary blow-up; exact large solutions; Hamilton–Jacobi–Bellman equations; stochastic optimal control; state constraints; viscosity solutions; operational research; inventory management; portfolio optimization
2020 Mathematics Subject Classification:
35J60 (Primary); 35B40, 35J75, 93E20, 49L25, 90B05, 91B70 (Secondary)
The final version of the article is published in journal (with volume, page numbers and DOI, when available).
1. Introduction
1.1. Historical Background and State of the Art
The study of boundary blow-up solutions (also known as large solutions) for elliptic equations has a celebrated history, dating back to the pioneering works of Bieberbach [6] and Rademacher [24] on the equation $\Delta u=f(u)$ with $f(u)=e^{u}$. These early investigations showed that the blow-up rate is determined by the interplay between the domain geometry and the nonlinearity structure. The modern theory was significantly advanced by Keller [14] and Osserman [22], who established the necessary and sufficient condition for existence:
$\int_{1}^{\infty}[F(s)]^{-1/2}\,ds<\infty,\quad F(s)=\int_{0}^{s}f(t)\,dt.$ (1.1)
This Keller–Osserman condition has since been generalized to a wide class of operators and settings; see [20, 7] for comprehensive reviews.
In subsequent decades, the theory was extended in multiple directions. Marcus and Véron [20] developed a comprehensive framework for uniqueness and asymptotic behavior of solutions with boundary blow-up, establishing connections between the growth rate of the nonlinearity and the geometry of the domain. Díaz and Letelier [11] studied explosive solutions of quasilinear elliptic equations, introducing techniques that allow for more general differential operators. Bandle and Giarrusso [3] initiated the study of boundary blow-up for equations with nonlinear gradient terms, demonstrating that the interaction between the absorption term and the gradient term generates qualitatively new phenomena.
A pivotal development in the theory was the introduction of equations involving gradient-dependent nonlinearities. In such equations, the structure
$\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u=f(x)$ (1.2)
requires fundamentally different techniques from the classical Keller–Osserman framework, since the gradient term introduces a first-order singularity that competes with the second-order diffusion. The interplay between these terms determines the asymptotic behavior of the solution near the boundary, giving rise to distinct asymptotic regimes.
A modern frontier in the field is the connection to stochastic optimal control, as pioneered by Lasry and Lions [17]. In their seminal 1989 paper, they demonstrated that boundary blow-up solutions to Hamilton–Jacobi–Bellman (HJB) equations serve as value functions for control problems with state constraints. This dual interpretation provides both physical intuition and rigorous regularity results. Specifically, if a controlled diffusion process is required to remain within a domain $\Omega$ for all time (a state constraint), then the associated value function satisfies an elliptic PDE with $u(x)\to+\infty$ as $x\to\partial\Omega$. The infinite boundary value acts as an “infinite penalty” that prevents the optimal trajectory from exiting the domain.
This stochastic control interpretation has been further developed in the context of ergodic control [4], mean-field games [18], and viscosity solutions [10]; see also the article [8] and the monographs by Fleming and Soner [12] and Bensoussan and Lions [5]. Recent advances by Zhang [27] in the context of Monge-Ampère equations have introduced high-precision analysis for singular weights, which we adapt here for the Laplacian setting.
From the perspective of Operational Research, the connection between elliptic PDEs and stochastic control has far-reaching implications. The value function of a state-constrained stochastic control problem encodes the optimal cost-to-go for a manager who must keep a stochastic system within safe operating limits. This formulation arises naturally in inventory management [23, 28], portfolio optimization under risk constraints [21], and production planning in manufacturing systems [8, 26].
1.2. Limitations of Existing Models
Despite extensive research since the pioneering works of Bieberbach [6] and Rademacher [24], existing models frequently face limitations that hinder their general applicability. We identify four principal limitations:
Limitation 1: Geometric restrictions. Many classical studies are restricted to radial domains (balls), where the geometry reduces to a single ordinary differential equation. While these simplifications allow for elegant closed-form solutions, they fail to capture the complexity of real-world scenarios where the underlying state space is not symmetric. In economic applications, the feasible operating region is rarely a ball; it is typically a polyhedron, an ellipsoid, or a more general convex set defined by multiple constraints.
Limitation 2: Non-singular coefficients. The majority of existing results assume that the coefficients $a(x)$ and $b(x)$ are bounded and bounded away from zero. This assumption is incompatible with economic models where the cost of corrective action or the discount rate varies significantly with the state. For instance, in inventory management, the cost of emergency replenishment increases dramatically as stock approaches zero, corresponding to a weight $b(x)$ that blows up at the boundary.
Limitation 3: Absence of gradient terms. The Keller–Osserman condition (1.1) applies to equations of the form $\Delta u=f(u)$, where there is no gradient dependence. The introduction of a gradient term $h(|\nabla u|)$ fundamentally changes the problem structure, and the classical theory does not directly apply. A new balance equation must be derived that accounts for the competition between the Laplacian and the gradient term.
Limitation 4: Gap between theory and practice. Most existing results provide existence theorems under abstract conditions that are difficult to translate into operational control laws for managers or engineers. There is a significant gap between purely analytical investigations and their practical implementation. The theory must provide not only existence and uniqueness, but also explicit asymptotic formulas, convexity guarantees, and numerical algorithms that can be implemented in practice.
1.3. The New General Framework
The motivation for this work is the construction of a general theory for the class of singular problems described by (1.2). Our framework addresses all four limitations identified above:
  1. (i)
    We work on general bounded strictly convex domains $\Omega\subset\mathbb{R}^{N}$, not restricting to radial geometries.
  2. (ii)
    We allow the weights $a(x)$ and $b(x)$ to exhibit prescribed singular behavior near $\partial\Omega$, with explicit power-law rates.
  3. (iii)
    We handle general strictly convex Hamiltonians $h$ with power-type growth $h(s)\sim s^{q}$ for $q\in(1,2]$, fully accounting for the competition between diffusion and gradient terms.
  4. (iv)
    We provide explicit asymptotic constants, convexity results, and a stochastic control interpretation that directly yields optimal feedback control laws.
Our approach is built on the principle of independence from concrete data meaning. We argue that the fundamental dynamics of “preventive correction” are governed by universal geometric and analytical laws. This allows the researcher to apply our results to any system where a state $X_{t}$ must be kept in $\Omega$ by a drift $\xi_{t}$, provided the cost functional mirrors the singular structures we define. In this sense, the article provides general, high-precision tools in stochastic optimization that remain valid as long as the structural assumptions are met.
The key innovation in our model lies in the identification of three distinct asymptotic regimes determined by the relationship between the gradient exponent $q$ and the weight singularity $\beta$:
  • Gradient-Dominant Regime ($q<\beta+2$): The solution blows up as $u(x)\sim C\,d(x)^{-\gamma}$ with $\gamma=(\beta-q+2)/(q-1)>0$.
  • Critical Logarithmic Regime ($q=\beta+2$): The blow-up is logarithmic, $u(x)\sim C\,\ln(1/d(x))$.
  • High-Order Gradient Regime ($q>\beta+2$): The gradient term dominates at all orders.
This classification, combined with the exact analytical constants we derive, provides a complete picture of the boundary behavior. The universal correction constant $\xi_{0}$ that we identify represents the precise intensity of the boundary penalty required to enforce the state constraint, and it depends only on the structural parameters $(q,\beta,b_{0},l_{0})$ of the problem.
1.4. Main Contributions and Organization
In this paper, we study the semilinear elliptic equation
$\left\{\begin{array}[]{ll}-\Delta u(x)+b(x)\,h(|\nabla u(x)|)+a(x)\,u(x)=f(x)&\text{for }x\in\Omega,\\ u(x)\rightarrow+\infty&\text{as }\mathrm{dist}(x,\partial\Omega)\rightarrow 0,\end{array}\right$. (1.3)
where $\Omega\subset\mathbb{R}^{N}$ is a bounded strictly convex domain, $h$ is a strictly convex function with power-like growth, and the weights $a(x)$, $b(x)$ satisfy specific singular growth conditions near $\partial\Omega$.
Our main contributions are:
  1. (1)
    Existence and uniqueness (Theorem 2.9): We prove existence and uniqueness for a class of semilinear equations on strictly convex (non-radial) domains with singular weights $a(x)$ and $b(x)$, via Perron’s method and explicit barrier construction.
  2. (2)
    Sharp boundary asymptotics (Theorem 2.10): We identify three distinct regimes (gradient-dominant, high-order, and logarithmic) and derive the exact analytical constants that govern the blow-up rate, mirroring the rigorous approach of Zhang [27].
  3. (3)
    Convexity inheritance (Theorem 2.11): Under the additional assumptions that $a$ and $b$ are convex and $f$ is concave, we prove that the strict convexity of the boundary is inherited by the solution’s Hessian, ensuring the stability and uniqueness of optimal control strategies.
  4. (4)
    Verification theorem (Theorem 6.9): We prove that the analytical solution is the value function of a state-constrained stochastic control problem and derive the optimal feedback control law explicitly.
  5. (5)
    Economic applications: We develop a comprehensive section on applications in Operational Research and Management Science, including real-world case studies and interdisciplinary connections with physics.
  6. (6)
    Numerical simulations for $N=1$ and $N=2$: We implement and validate the theory numerically in both one and two spatial dimensions, with detailed interpretation of the results.
1.5. Methodology: The Calculus of Singular Balancing
Our methodology proceeds in four phases, designed to bridge the gap between abstract analysis and practical economic applicability:
Phase 1: Singularity Balance Identification.:
We analyze the HJB equation as a competition between Diffusion ($\Delta u$) and Correction ($b\,h(|\nabla u|)$). By substituting the ansatz $u\sim d^{-\gamma}$, we find the value of $\gamma$ that matches the singular orders. This phase identifies the three asymptotic regimes.
Phase 2: Barrier Construction.:
We use the concave defining function $v(x)$ to construct global sub- and supersolutions that determine the precise boundary blow-up rate. This phase ensures that local boundary estimates propagate correctly to the interior of the domain.
Phase 3: Stochastic Embedding.:
The PDE solution is mapped to a stochastic control problem. The blow-up $u\to\infty$ is interpreted as an impenetrable “cost wall” that enforces the safe-zone boundaries. The Legendre transform mediates the connection between the Hamiltonian and the running cost.
Phase 4: Economic Variable Mapping.:
Mathematical constants are mapped to economic parameters: $a(x)$ to state-dependent discount rates, $b(x)$ to adjustment stiffness coefficients, and $h$ to convex adjustment cost functions. This phase enables direct application to operational research problems.
1.6. Structure of the Paper
The remainder of the paper is organized as follows. Section 2 presents the mathematical framework and all structural assumptions. Section 3 develops the geometric preliminaries and barrier construction. Section 4 provides the complete proof of existence, uniqueness, and convexity (Theorem 2.9 and Theorem 2.11). Section 5 details the proof of sharp boundary asymptotics (Theorem 2.10). Section 6 develops the stochastic control model, the verification theorem, and the full derivation of the Hamilton–Jacobi–Bellman equation. Section 7 presents applications in Operational Research and Management Science, including case studies and interdisciplinary connections with physics. Section 8 describes the monotone iterative scheme. Section 9 presents numerical simulations for both the $N=1$ and $N=2$ cases with detailed interpretation. Section 10 concludes the work. Appendix A and Appendix B contain the complete Python implementation codes for the one-dimensional and two-dimensional solvers, respectively.
2. Mathematical Framework and Assumptions
Throughout this paper, we adopt the following structural and geometric assumptions. These assumptions are designed to be as general as possible while ensuring that the main results (existence, uniqueness, asymptotics, convexity, and stochastic representation) hold simultaneously.
Assumption 2.1 (Domain).
Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 1$) be a bounded open set. The boundary $\partial\Omega$ is of class $C^{2,\alpha}$ for some $\alpha\in(0,1)$. We assume $\Omega$ is strictly convex: there exists a positive constant $\kappa_{0}>0$ such that for every $x\in\partial\Omega$, the principal curvatures $\kappa_{i}(x)$ satisfy:
$\min_{i=1,\dots,N-1}\kappa_{i}(x)\geq\kappa_{0}.$ (2.1)
Remark 2.2 (Geometric significance of strict convexity).
Strict convexity is a natural assumption in both mathematics and economics. In economic dynamics, the domain $\Omega$ represents the set of feasible states (e.g., inventory levels, portfolio weights, production capacities). Strict convexity ensures that any convex combination of two feasible states is also feasible, which is the mathematical formulation of the principle that “diversification reduces risk.” The curvature bound (2.1) ensures that the boundary has no flat segments or corners, which would create singular shocks in the optimal control strategy. In physical terms, the curvature controls the rate at which the distance function $d(x)=\mathrm{dist}(x,\partial\Omega)$ becomes non-smooth, and the lower bound $\kappa_{0}$ guarantees that the tubular neighborhood of $\partial\Omega$ where $d(x)$ is $C^{2}$ has width at least $1/\kappa_{0}$.
Assumption 2.3 (Source term).
$f:\overline{\Omega}\to[0,\infty)$ is a concave function belonging to $C^{k,\alpha}(\overline{\Omega})$ for some $k\geq 1$ and $\alpha\in(0,1)$.
Assumption 2.4 (Hamiltonian).
The function $h:[0,\infty)\to[0,\infty)$ satisfies:
  1. (H1)
    $h\in C^{2}((0,\infty))\cap C([0,\infty))$ with $h(0)=0$;
  2. (H2)
    $h$ is strictly increasing for all $s>0$ and strictly convex;
  3. (H3)
    There exist constants $l_{1},l_{2}>0$ and $q\in(1,2]$ such that
    $l_{1}\,s^{q}\leq h(s)\leq l_{2}\,s^{q}\quad\text{for all }s\geq 0.$ (2.2)
Remark 2.5 (Economic interpretation of the Hamiltonian).
The Hamiltonian $h$ represents the cost of corrective action per unit of adjustment intensity. The strict convexity (H2) captures the economic principle of increasing marginal costs: the cost of doubling the adjustment rate is more than double the original cost. The power-type growth (H3) with exponent $q\in(1,2]$ encompasses the quadratic case $h(s)=s^{2}$ (commonly used in linear-quadratic control) as well as sub-quadratic costs that arise in models with capacity constraints. The canonical example is $h(s)=s^{q}/q$ for $q\in(1,2]$.
Assumption 2.6 (Weights).
The weights $a,b\in C^{\infty}(\Omega)\cap C^{k,\alpha}(\Omega)$ for $k\geq 1$, are strictly positive in $\Omega$ and satisfy:
  1. (W1)
    There exists a constant $\alpha>-2$ and $a_{1},a_{2}>0$ such that
    $a_{1}\,d(x)^{\alpha}\leq a(x)\leq a_{2}\,d(x)^{\alpha}$ (2.3)
    near $\partial\Omega$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. We further assume that $a(x)$ is positive convex in $\Omega$.
  2. (W2)
    There exists a constant $\beta\in\mathbb{R}_{+}=[0,+\infty)$ and limits $b_{1},b_{2}>0$ such that
    $\liminf_{d(x)\to 0}\frac{b(x)}{d(x)^{\beta}}=b_{1}\leq\limsup_{d(x)\to 0}\frac{b(x)}{d(x)^{\beta}}=b_{2}.$ (2.4)
    We assume that $b(x)$ is positive convex in $\Omega$.
Remark 2.7 (Principle of Independence from Data Meaning).
Assumptions 2.12.6 do not assume a specific physical meaning for $x$. This abstraction is intentional: we seek a theory where the boundary dynamics are determined by the category of the singularity (characterized by the exponents $\alpha$, $\beta$, $q$), not by the specific data values. The same equation may describe inventory levels in a warehouse, concentrations in a chemical reactor, or asset allocations in a portfolio, and our results apply identically to all these scenarios.
This structure on the weights mirrors the rigorous bounds employed by Zhang [27] in the context of Monge-Ampère equations with weights, enabling us to derive precise trace limits for our solutions.
Remark 2.8 (Applicability of the Lasry–Lions framework).
The assumptions (H1)–(H3) for the Hamiltonian $h$ together with the strict convexity of $\Omega$ (Assumption 2.1) ensure that the fundamental requirements of the Lasry–Lions framework [17] are fully satisfied. In particular, the strict convexity of $h$ and its power-law growth $h(s)\sim s^{q}$ for $q>1$, combined with the positivity and singular growth of the reaction coefficient $a(x)$, guarantee that the boundary blow-up solution $u$ in (1.3) arises as the value function of a well-posed state-constrained stochastic control problem. Thus, the analytical problem considered in this work is consistently embedded within the classical stochastic paradigm introduced in [17].
2.1. Statement of Main Results
We now state our main contributions precisely.
Theorem 2.9 (Existence and Uniqueness).
Under Assumptions 2.12.6, the HJB equation (1.3) admits a unique classical solution $u\in C^{2}(\Omega)$ such that $u(x)\to\infty$ as $d(x)\to 0$. In the Gradient-Dominant regime ($q<\beta+2$), the solution satisfies:
$C_{-}\,d(x)^{-\gamma}\leq u(x)\leq C_{+}\,d(x)^{-\gamma}\quad\text{near }\partial\Omega,\quad\gamma=\frac{\beta-q+2}{q-1},$ (2.5)
where $C_{\pm}$ are explicit barrier constants defined in Lemma 3.2.
Theorem 2.10 (Exact Sharp Boundary Asymptotics).
Let the conditions of Theorem 2.9 hold, and assume further that
$\lim_{s\to\infty}h(s)/s^{q}=l_{0}\in[l_{1},l_{2}].$
Then the solution $u(x)$ satisfies the sharp boundary limit:
$\xi_{1}\leq\liminf_{d(x)\to 0}u(x)\,(d(x))^{\gamma}\leq\limsup_{d(x)\to 0}u(x)\,(d(x))^{\gamma}\leq\xi_{2},$ (2.6)
where the analytic limit bounds $\xi_{1},\xi_{2}$ are defined as:
$\xi_{1}=\left(\frac{\gamma(\gamma+1)}{b_{2}\,l_{2}\,\gamma^{q}}\right)^{\frac{1}{q-1}},\quad\xi_{2}=\left(\frac{\gamma(\gamma+1)}{b_{1}\,l_{1}\,\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (2.7)
If the limits
$b_{0}=\lim_{d(x)\to 0}b(x)\,d(x)^{-\beta}\quad\text{and}\quad l_{0}=\lim_{s\to\infty}h(s)\,s^{-q}$
exist, then:
$\lim_{d(x)\to 0}u(x)\,(d(x))^{\gamma}=\xi_{0}=\left(\frac{\gamma(\gamma+1)}{b_{0}\,l_{0}\,\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (2.8)
The constant $\xi_{0}$ is the Universal Correction Constant. It represents the precise intensity of the boundary penalty required to enforce the state constraint. Its explicit dependence on $(q,\beta,b_{0},l_{0})$ provides a quantitative tool for calibrating control strategies.
Theorem 2.11 (Convexity Inheritance).
Under Assumptions 2.12.6, if the source term $f(x)$ is concave and the reaction weights $a,b$ are convex in $\Omega$, then the solution $u$ from Theorem 2.9 is strictly convex throughout $\Omega$: $D^{2}u(x)>0$ for all $x\in\Omega$.
Strict convexity is indispensable for practical control: it ensures that the optimal decision for any state $x$ is unique and can be found via simple gradient-descent algorithms. It also guarantees that the running cost functional is jointly convex, which is a prerequisite for the application of duality theory in optimization.
3. Geometric Preliminaries and Barrier Construction
3.1. The Defining Function
The strict convexity of $\Omega$ (Assumption 2.1) implies the existence of a smooth strictly convex defining function. Following the approach of Gilbarg and Trudinger [13] and Lazer and McKenna [19], we establish:
Lemma 3.1 (Convex defining function).
Under Assumption 2.1, there exists $\varphi\in C^{2}(\overline{\Omega})$ satisfying:
  1. (1)
    $\varphi(x)<0$ for all $x\in\Omega$;
  2. (2)
    $\varphi(x)=0$ for all $x\in\partial\Omega$;
  3. (3)
    $|\nabla\varphi(x)|=1$ for all $x\in\partial\Omega$;
  4. (4)
    $D^{2}\varphi(x)\geq\kappa_{0}\,I$ for all $x\in\overline{\Omega}$, where $\kappa_{0}>0$ is the minimum principal curvature of $\partial\Omega$ from (2.1).
Proof.
The proof proceeds in two steps.
Step 1: Construction near the boundary. For $x$ sufficiently close to $\partial\Omega$, define $\varphi(x)=-d(x)$, where $d(x)=\mathrm{dist}(x,\partial\Omega)$. The signed distance function is $C^{2}$ in a tubular neighborhood $\mathcal{T}_{\delta}=\{x\in\Omega:d(x)<\delta\}$ when the boundary is $C^{2,\alpha}$ (see [13, Chapter 14]). The width of this neighborhood is at least $\delta=1/\kappa_{0}$, where $\kappa_{0}$ is the maximum of the principal curvatures.
The Hessian of d-d at a point $x\in\mathcal{T}_{\delta}$ is related to the second fundamental form of the level set $\{d=\text{const}\}$. By the strict convexity of $\Omega$ (Assumption 2.1), the principal curvatures of these level sets are positive, which gives $D^{2}(-d)(x)\geq\kappa_{0}\,I$ in $\mathcal{T}_{\delta}$.
Step 2: Extension to the interior. The function $\varphi=-d$ can be extended to all of $\overline{\Omega}$ while preserving strict convexity. This is achieved by considering the convex envelope of d-d in the interior region $\Omega\setminus\mathcal{T}_{\delta/2}$ and smoothing via convolution. The detailed construction is given in [13, Chapter 14]. ∎
Let $v(x)=-\varphi(x)$. Then $v>0$ in $\Omega$, $v=0$ on $\partial\Omega$, and $v$ is strictly concave ($D^{2}v<0$). By construction, $v(x)=d(x)$ identically in a tubular neighborhood of $\partial\Omega$, where $|\nabla v(x)|=1$ holds. These properties of $v$ are used extensively in the barrier construction that follows.
3.2. Explicit Barrier Construction
We now construct explicit sub- and supersolutions that determine the precise boundary blow-up rate. This construction is the technical core of the paper.
Lemma 3.2 (Explicit Barriers).
Let $W(x)=C\,v(x)^{-\gamma}$, where $v(x)=-\varphi(x)$ is the concave defining function from Lemma 3.1. Define the elliptic operator
$\mathcal{L}u=-\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u-f(x).$ (3.1)
Under Assumptions 2.12.6, the following hold:
  1. (1)
    Gradient-Dominant Case ($q<\beta+2$): The main singularity is governed by the balance between the diffusion $\Delta u$ and the gradient term $b(x)\,h(|\nabla u|)$. Setting $\gamma=(\beta-q+2)/(q-1)>0$ and the critical constant
    $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}\,l_{1}\,\gamma^{q}}\right)^{\frac{1}{q-1}},$ (3.2)
    there exist $C_{+}>C^{*}$ and $0<C_{-}<C^{*}$ such that $\overline{u}(x)=C_{+}\,v(x)^{-\gamma}$ is a supersolution and $\underline{u}(x)=C_{-}\,v(x)^{-\gamma}$ is a subsolution in a neighborhood of $\partial\Omega$.
  2. (2)
    High-Order Gradient Case ($q>\beta+2$): The gradient term grows faster than the Laplacian as $x\to\partial\Omega$. For any $\gamma>0$, the gradient term dominates, ensuring that $\mathcal{L}(C\,v^{-\gamma})>0$ for small $v$ and any $C>0$.
  3. (3)
    Critical Logarithmic Case ($q=\beta+2$): The singular growth is logarithmic, $u(x)\sim C\,\ln(1/v(x))$. There exist $C_{+}>C_{-}^{*}>0$ and $0<C_{-}<C_{-}^{*}$ providing local barriers near the boundary.
Proof.
We provide detailed computations for each case, using all hypotheses explicitly.
Step 1: Differential calculus for $W=C\,v^{-\gamma}$.
We compute the gradient, its norm, and the Laplacian of $W$, using only the chain rule and the properties of $v$ from Lemma 3.1:
$\nabla W$ $=-C\,\gamma\,v^{-\gamma-1}\,\nabla v,$ (3.3)
$|\nabla W|$ $=C\,\gamma\,v^{-\gamma-1}\,|\nabla v|,$ (3.4)
$\Delta W$ $=-C\,\gamma\left[-(\gamma+1)\,v^{-\gamma-2}\,|\nabla v|^{2}+v^{-\gamma-1}\,\Delta v\right]$
$=C\,\gamma\,(\gamma+1)\,v^{-\gamma-2}\,|\nabla v|^{2}-C\,\gamma\,v^{-\gamma-1}\,\Delta v.$ (3.5)
For the Hessian of $W$, we compute:
$D^{2}W=C\,\gamma\,(\gamma+1)\,v^{-\gamma-2}\,\nabla v\otimes\nabla v-C\,\gamma\,v^{-\gamma-1}\,D^{2}v.$ (3.6)
Step 2: Substitution into the operator $\mathcal{L}$.
Substituting (3.3)–(3.5) into $\mathcal{L}W$ defined in (3.1):
$\mathcal{L}W$ $=-\Delta W+b(x)\,h(|\nabla W|)+a(x)\,W-f(x)$
$=-C\,\gamma\,(\gamma+1)\,v^{-\gamma-2}\,|\nabla v|^{2}+C\,\gamma\,v^{-\gamma-1}\,\Delta v$
$\quad+b(x)\,h\bigl(C\,\gamma\,v^{-\gamma-1}\,|\nabla v|\bigr)+a(x)\,C\,v^{-\gamma}-f(x).$ (3.7)
Step 3: Asymptotic analysis near $\partial\Omega$ for Case 1.
We now use the weight hypotheses (W1)–(W2) from Assumption 2.6 and the Hamiltonian bounds (H3) from Assumption 2.4. Near $\partial\Omega$, the properties of $v$ from Lemma 3.1 give $v(x)=d(x)$ and $|\nabla v(x)|=1$.
For any $\varepsilon>0$, by (2.4), there exists $\delta>0$ such that for all $x\in\Omega$ with $v(x)<\delta$:
$(b_{1}-\varepsilon)\,v(x)^{\beta}\leq b(x)\leq(b_{2}+\varepsilon)\,v(x)^{\beta}.$ (3.8)
Using the lower bound in (H3), the gradient term satisfies:
$b(x)\,h(C\,\gamma\,v^{-\gamma-1}\,|\nabla v|)$ $\geq(b_{1}-\varepsilon)\,v^{\beta}\cdot l_{1}\,(C\,\gamma)^{q}\,v^{-q(\gamma+1)}$
$=(b_{1}-\varepsilon)\,l_{1}\,(C\,\gamma)^{q}\,v^{\beta-q(\gamma+1)}.$ (3.9)
The leading singular terms in $\mathcal{L}W$ are:
$\mathcal{L}W\geq v^{-\gamma-2}\left[(b_{1}-\varepsilon)\,l_{1}\,(C\,\gamma)^{q}\,v^{\beta-q(\gamma+1)+\gamma+2}-C\,\gamma\,(\gamma+1)\right]+\mathcal{R}(x),$ (3.10)
where $\mathcal{R}(x)$ denotes terms of strictly lower singular order relative to $v^{-\gamma-2}$, specifically:
$\mathcal{R}(x)=C\,\gamma\,v^{-\gamma-1}\,\Delta v+a(x)\,C\,v^{-\gamma}-f(x).$ (3.11)
Step 4: Balance of leading singularities.
We choose $\gamma$ to perform an exact balance of the leading singularities by requiring the exponent of $v$ inside the brackets in (3.10) to vanish:
$\beta-q(\gamma+1)+\gamma+2=0\iff\gamma(1-q)=\beta-q+2\iff\gamma=\frac{\beta-q+2}{q-1}.$ (3.12)
Note that $\gamma>0$ if and only if $\beta-q+2>0$, i.e., $q<\beta+2$. This is the Gradient-Dominant regime.
With this choice of $\gamma$, the expression in brackets becomes:
$P(C)=(b_{1}-\varepsilon)\,l_{1}\,\gamma^{q}\,C^{q}-\gamma\,(\gamma+1)\,C.$ (3.13)
Since $q>1$:
  • $P(C)<0$ for small $C>0$ (the linear term $\gamma(\gamma+1)C$ dominates);
  • $P(C)>0$ for large $C$ (the power term $C^{q}$ dominates);
  • $P(C^{*})=0$ at the critical constant $C^{*}=\left(\frac{\gamma(\gamma+1)}{b_{1}\,l_{1}\,\gamma^{q}}\right)^{1/(q-1)}$.
Therefore, choosing $C_{+}>C^{*}$ ensures $\mathcal{L}(C_{+}\,v^{-\gamma})>0$ (supersolution), and choosing $C_{-}<C^{*}$ ensures $\mathcal{L}(C_{-}\,v^{-\gamma})<0$ (subsolution) near $\partial\Omega$.
Step 5: Verification of the subordinate terms.
The reaction term, using (W1) from Assumption 2.6:
$a(x)\,C\,v^{-\gamma}\leq a_{2}\,C\,v^{\alpha-\gamma}.$ (3.14)
Since $\alpha>-2$ by (W1), we have $\alpha-\gamma>-2-\gamma=-\gamma-2$, so the reaction term is of order $v^{\alpha-\gamma}$, which is strictly lower than the leading order $v^{-\gamma-2}$. Therefore, after multiplying by $v^{\gamma+2}$, this term vanishes in the limit $v\to 0$.
The term $C\,\gamma\,v^{-\gamma-1}\,\Delta v$ is $O(v^{-\gamma-1})$, which is also lower order than $v^{-\gamma-2}$.
The source term $f(x)$ is bounded on $\overline{\Omega}$ by Assumption 2.3, hence $f(x)=O(1)=o(v^{-\gamma-2})$.
Step 6: Case 2 — High-order gradient.
When $q>\beta+2$, equation (3.12) gives $\gamma<0$ (the numerator $\beta-q+2$ is negative). Thus, for any chosen $\gamma>0$, the exponent $\beta-q(\gamma+1)+\gamma+2<0$, meaning $v^{\beta-q(\gamma+1)}\cdot v^{\gamma+2}$ has a negative exponent. The gradient term is therefore more singular than the Laplacian term, and $\mathcal{L}W>0$ near $\partial\Omega$ for any $C>0$.
Step 7: Case 3 — Critical logarithmic.
When $q=\beta+2$, let $W=C\,\ln(1/v)=-C\,\ln v$. Then:
$\nabla W$ $=-C\,v^{-1}\,\nabla v,$ (3.15)
$|\nabla W|$ $=C\,v^{-1}\,|\nabla v|,$ (3.16)
$\Delta W$ $=C\,v^{-2}\,|\nabla v|^{2}-C\,v^{-1}\,\Delta v.$ (3.17)
Substituting into $\mathcal{L}W$ and using $|\nabla v|=1$, $b(x)\geq(b_{1}-\varepsilon)\,v^{\beta}$, $h(s)\geq l_{1}\,s^{q}$:
$\mathcal{L}W$ $\geq-C\,v^{-2}+C\,v^{-1}\,\Delta v+(b_{1}-\varepsilon)\,v^{\beta}\,l_{1}\,(C\,v^{-1})^{q}+a_{1}\,v^{\alpha}\,C\,\ln(1/v)-f$
$=v^{-2}\left[(b_{1}-\varepsilon)\,l_{1}\,C^{q}\,v^{\beta-q+2}-C\right]+\mathcal{R}_{L}(x),$ (3.18)
where $\mathcal{R}_{L}(x)$ denotes terms of strictly lower singular order. Since $\beta-q+2=0$ in the critical case, we obtain:
$\liminf_{x\to\partial\Omega}\left(v^{2}\,\mathcal{L}W\right)\geq(b_{1}-\varepsilon)\,l_{1}\,C^{q}-C.$ (3.19)
The same argument as in Step 4 shows that barriers exist with the appropriate choice of $C_{\pm}$. ∎
Remark 3.3 (Role of strict convexity).
The strict convexity of $\Omega$ is essential in two ways: (i) it guarantees the existence of a smooth, strictly concave defining function $v$ (Lemma 3.1), ensuring $D^{2}v<0$; (ii) it ensures that $|\nabla v|$ is bounded away from zero near $\partial\Omega$, which is necessary for the gradient term to provide the required singular behavior. Without strict convexity, the gradient could vanish at flat boundary points, destroying the singularity balance.
4. Existence, Uniqueness, and Convexity
We now establish the main existence and uniqueness result using Perron’s method combined with the barrier functions from Section 3.
Lemma 4.1 (Comparison principle).
Let $\Omega^{\prime}\Subset\Omega$ be a smooth subdomain. Suppose $u,w\in C^{2}(\Omega^{\prime})\cap C(\overline{\Omega^{\prime}})$ satisfy
$\mathcal{L}u$ $\leq\mathcal{L}w\quad\text{in }\Omega^{\prime},$
$u$ $\leq w\quad\text{on }\partial\Omega^{\prime}.$ (4.1)
Then $u\leq w$ in $\Omega^{\prime}$.
Proof.
Suppose for contradiction that
$\max_{\overline{\Omega^{\prime}}}(u-w)=u(x_{0})-w(x_{0})>0$
for some $x_{0}\in\Omega^{\prime}$. The maximum cannot be achieved on $\partial\Omega^{\prime}$ by the boundary hypothesis (4.1). At the interior maximum point $x_{0}$:
$\nabla u(x_{0})$ $=\nabla w(x_{0}),$ (4.2)
$\Delta u(x_{0})$ $\leq\Delta w(x_{0}).$ (4.3)
Using (4.2), the Hamiltonian terms coincide: $h(|\nabla u(x_{0})|)=h(|\nabla w(x_{0})|)$. Therefore:
$0$ $\geq\mathcal{L}u(x_{0})-\mathcal{L}w(x_{0})$
$=-\Delta u(x_{0})+\Delta w(x_{0})+b(x_{0})\bigl[h(|\nabla u(x_{0})|)-h(|\nabla w(x_{0})|)\bigr]$
$\quad+a(x_{0})\bigl[u(x_{0})-w(x_{0})\bigr]$
$\geq a(x_{0})\bigl[u(x_{0})-w(x_{0})\bigr]>0,$
where we used (4.3) for the first term, (4.2) for the vanishing of the Hamiltonian difference, and $a(x_{0})>0$ from Assumption 2.6. This is a contradiction. ∎
Proof of Theorem 2.9.
The proof proceeds in five steps.
Step 1: Construction of global barriers.
Let $\overline{u}(x)=C_{+}\,v(x)^{-\gamma}$ and $\underline{u}(x)=C_{-}\,v(x)^{-\gamma}$ be the super- and subsolutions from Lemma 3.2. By construction, these satisfy
$\mathcal{L}\overline{u}\geq 0,\quad\mathcal{L}\underline{u}\leq 0\quad\text{in }\{x\in\Omega:d(x)<\delta_{0}\}$
for some $\delta_{0}>0$ determined by the asymptotic estimates in Lemma 3.2.
For the interior region $\Omega_{\delta_{0}}=\{x\in\Omega:d(x)\geq\delta_{0}\}$, we solve the Dirichlet problem
$\left\{\begin{array}[]{ll}\mathcal{L}u_{0}=0&\text{in }\Omega_{\delta_{0}},\\ u_{0}=\overline{u}&\text{on }\partial\Omega_{\delta_{0}}.\end{array}\right$. (4.4)
By standard quasilinear elliptic theory [13], this problem has a unique classical solution. By adjusting $C_{+}$ if necessary, we may assume $\overline{u}\geq u_{0}$ in $\Omega_{\delta_{0}}$.
Define the global supersolution as
$\overline{U}(x)=\begin{cases}\overline{u}(x)&\text{if }d(x)<\delta_{0},\\ \max\{\overline{u}(x),\,u_{0}(x)\}&\text{if }d(x)\geq\delta_{0}.\end{cases}$
Similarly construct a global subsolution $\underline{U}$ with $\underline{U}\leq\overline{U}$ throughout $\Omega$.
Step 2: Perron’s method.
Define the Perron class
$\mathcal{S}=\bigl\{w\in C(\Omega):w\text{ is a subsolution},\;\underline{U}\leq w\leq\overline{U}\bigr\}.$
The class $\mathcal{S}$ is nonempty since $\underline{U}\in\mathcal{S}$. Define the Perron solution
$u(x)=\sup_{w\in\mathcal{S}}w(x).$ (4.5)
Step 3: $u$ is a classical solution.
Let $x_{0}\in\Omega$ and let $B_{r}(x_{0})\Subset\Omega$ be a ball. For any $w\in\mathcal{S}$, let $\tilde{w}$ be the solution to
$\left\{\begin{array}[]{ll}\mathcal{L}\tilde{w}=0&\text{in }B_{r}(x_{0}),\\ \tilde{w}=w&\text{on }\partial B_{r}(x_{0}).\end{array}\right$.
By the comparison principle (Lemma 4.1), $w\leq\tilde{w}\leq\overline{U}$ in $B_{r}(x_{0})$. The “bump” construction shows that the function
$\hat{w}(x)=\begin{cases}\tilde{w}(x)&x\in B_{r}(x_{0}),\\ w(x)&x\in\Omega\setminus B_{r}(x_{0})\end{cases}$
is also a subsolution, hence $\hat{w}\in\mathcal{S}$.
Taking the supremum over all such modifications, standard arguments [10] show that $u$ is a viscosity solution. By interior regularity for uniformly elliptic equations with Hölder continuous coefficients, $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$, and $u$ is a classical solution. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13].
Step 4: Uniqueness.
Suppose $u_{1}$ and $u_{2}$ are two solutions satisfying (2.5). Consider $w=u_{1}-u_{2}$. Then $w$ satisfies a linearized equation of the form
$\Delta w+c(x)\cdot\nabla w+e(x)\,w=0,$ (4.6)
where the coefficients $c(x)$ and $e(x)$ are obtained from the mean value theorem applied to $h(|\nabla u|)$ and a(x)ua(x)\,u:
$c(x)$ $=b(x)\,\frac{h(|\nabla u_{1}|)-h(|\nabla u_{2}|)}{|\nabla u_{1}|-|\nabla u_{2}|}\,\frac{\nabla(|\nabla u_{1}|-|\nabla u_{2}|)}{|\nabla u_{1}|-|\nabla u_{2}|},$
$e(x)$ $=a(x).$
Since $w=u_{1}-u_{2}=o(d^{-\gamma})$ as $d\to 0$ (because both $u_{1}$ and $u_{2}$ satisfy the same sharp asymptotics (2.5)), the function $w$ grows slower than the barriers near the boundary.
For any $\epsilon>0$, define $\Omega_{\epsilon}=\{x\in\Omega:d(x)>\epsilon\}$. On $\partial\Omega_{\epsilon}$, both $u_{1}$ and $u_{2}$ are finite. The maximum principle applied to (4.6) on $\Omega_{\epsilon}$ gives:
$\sup_{\Omega_{\epsilon}}|w|\leq\sup_{\partial\Omega_{\epsilon}}|w|.$
Since $|w|=|u_{1}-u_{2}|=o(d^{-\gamma})$ on $\partial\Omega_{\epsilon}$ (where $d=\epsilon$), and the precise asymptotic matching forces $u_{1}/u_{2}\to 1$ as $d\to 0$, we conclude that $w\equiv 0$ by taking $\epsilon\to 0$.
Step 5: Interior regularity.
The solution $u$ defined by (4.5) belongs to $C^{2,\alpha}(\Omega)$ for some $\alpha\in(0,1)$. If additionally $f\in C^{k,\alpha}(\overline{\Omega})$ and $a,b\in C^{k,\alpha}(\Omega)$ for $k\geq 1$, then $u\in C_{\mathrm{loc}}^{k+2,\alpha}(\Omega)$. This follows from standard Schauder estimates for quasilinear elliptic equations; see [13, Chapter 13]. ∎
Proof of Theorem 2.11.
The proof of strict convexity proceeds in four steps, using the microscopic convexity principle following Kennington [15] and Korevaar [16].
Step 1: Setup.
Let $u$ be the solution from Theorem 2.9. Define the minimum eigenvalue function
$\lambda(x)=\min_{|\xi|=1}D^{2}u(x)\,\xi\cdot\xi.$ (4.7)
We aim to show $\lambda(x)>0$ for all $x\in\Omega$.
Step 2: Differential inequality.
Suppose $\lambda$ achieves a non-positive minimum at some $x_{0}\in\Omega$. At $x_{0}$, let $\xi_{0}$ be the corresponding eigenvector. Differentiating the equation $\mathcal{L}u=0$ twice in the direction $\xi_{0}$, and using the fact that $h$ is strictly convex (Assumption 2.4, condition (H2)), we obtain:
$\Delta\lambda+c(x)\cdot\nabla\lambda+e(x)\,\lambda\leq\mathcal{Q}(x,u,\nabla u)\quad\text{at }x_{0},$ (4.8)
where $c$ and $e$ are bounded functions, and the quadratic form $\mathcal{Q}$ is given by:
$\mathcal{Q}=-\bigl[b^{\prime\prime}(x)\,h(|\nabla u|)+a^{\prime\prime}(x)\,u-f^{\prime\prime}(x)\bigr],$ (4.9)
where $b^{\prime\prime}$, $a^{\prime\prime}$, $f^{\prime\prime}$ denote second directional derivatives in the direction $\xi_{0}$.
Under the hypotheses that $a(x)$ and $b(x)$ are convex (Assumption 2.6) and $f(x)$ is concave (Assumption 2.3), we have $a^{\prime\prime}\geq 0$, $b^{\prime\prime}\geq 0$, and $f^{\prime\prime}\leq 0$. Since $h\geq 0$, $u>0$, and $f\geq 0$, we conclude:
$\mathcal{Q}(x,u,\nabla u)\leq 0.$ (4.10)
Step 3: Boundary behavior.
Near $\partial\Omega$, the solution $u$ is well-approximated by the barrier $W=C\,v^{-\gamma}$. The Hessian of $W$, computed in (3.6), is:
$D^{2}W=C\,\gamma\,(\gamma+1)\,v^{-\gamma-2}\,\nabla v\otimes\nabla v-C\,\gamma\,v^{-\gamma-1}\,D^{2}v.$
Since $v$ is strictly concave ($D^{2}v<0$ by Lemma 3.1), the term $C\,\gamma\,v^{-\gamma-1}\,D^{2}v$ is positive definite. The rank-one term $\nabla v\otimes\nabla v$ is positive semi-definite. Therefore, $D^{2}W>0$ near $\partial\Omega$, which by comparison forces $D^{2}u>0$ near the boundary.
Step 4: Strong maximum principle.
Suppose $\lambda(x_{0})=\min_{\overline{\Omega}}\lambda\leq 0$ for some $x_{0}\in\Omega$. By Step 3, $x_{0}$ cannot be near $\partial\Omega$. By the differential inequality (4.8) with the favorable sign (4.10), and the strong maximum principle for linear elliptic equations, either $\lambda\equiv\lambda(x_{0})$ (constant) or the minimum cannot be achieved in the interior.
If $\lambda$ were constant and non-positive, this would contradict the positive definiteness of $D^{2}u$ near $\partial\Omega$ established in Step 3. Therefore, $\lambda>0$ throughout $\Omega$, which means $D^{2}u(x)>0$ for all $x\in\Omega$. ∎
5. Sharp Boundary Asymptotics
Proof of Theorem 2.10.
The derivation of the limit constants $\xi_{1},\xi_{2}$ relies on a precise asymptotic balance in the equation $\mathcal{L}u=0$ as $d(x)\to 0$. Let $u$ be the unique solution from Theorem 2.9 and assume the gradient-dominant case $q<\beta+2$.
Step 1: Singular order balance.
As established in Section 3, for a blow-up rate $\gamma$, the individual terms in the equation scale as follows near $\partial\Omega$:
$\Delta u$ $\sim C\,\gamma\,(\gamma+1)\,d(x)^{-\gamma-2},$
$b(x)\,h(|\nabla u|)$ $\sim b(x)\,l_{0}\,(C\,\gamma)^{q}\,d(x)^{-q(\gamma+1)}$
$\sim b_{0}\,l_{0}\,(C\,\gamma)^{q}\,d(x)^{\beta-q(\gamma+1)},$
$a(x)\,u$ $\sim a_{0}\,C\,d(x)^{\alpha-\gamma},$
$f(x)$ $=O(1).$
Equating the singular orders of the two leading terms (Laplacian and weighted gradient):
$\gamma-2=\beta-q\gamma-q\implies\gamma(q-1)=\beta-q+2\implies\gamma=\frac{\beta-q+2}{q-1}.$ (5.1)
The reaction term a(x)u(x)a(x)\,u(x) behaves like $d(x)^{\alpha-\gamma}$. Since $\alpha>-2$ by Assumption 2.6, we have $\alpha-\gamma>-2-\gamma=-(\gamma+2)$. This confirms that the reaction term is of lower singular order than the leading terms $d(x)^{-\gamma-2}$ and vanishes after multiplying the equation by $d(x)^{\gamma+2}$ and taking the limit $d\to 0$.
Step 2: Constant identification.
Multiply the equation
$\Delta u+b(x)\,h(|\nabla u|)+a(x)\,u=f(x)$
by $d(x)^{\gamma+2}$. As $x\to\partial\Omega$, let $\xi=\lim_{d\to 0}u(x)\,d(x)^{\gamma}$ (assuming the limit exists). Then $u(x)\sim\xi\,d(x)^{-\gamma}$, and:
$-\Delta u$ $\sim-\xi\,\gamma\,(\gamma+1)\,d(x)^{-\gamma-2},$ (5.2)
$b(x)\,h(|\nabla u|)$ $\sim b(x)\,l_{0}\,(\xi\,\gamma)^{q}\,d(x)^{-q(\gamma+1)}.$ (5.3)
After multiplying by $d^{\gamma+2}$ and using (5.1), the equation becomes in the limit:
$\xi\,\gamma\,(\gamma+1)+b_{i}\,l_{i}\,(\xi\,\gamma)^{q}=0,$ (5.4)
where the subscript $i$ indicates whether we use the liminf or limsup of $b(x)\,d(x)^{-\beta}$ and $h(s)\,s^{-q}$.
Solving (5.4) for $\xi$:
$b_{i}\,l_{i}\,\gamma^{q}\,\xi^{q}=\gamma\,(\gamma+1)\,\xi\implies\xi^{q-1}=\frac{\gamma\,(\gamma+1)}{b_{i}\,l_{i}\,\gamma^{q}}\implies\xi=\left(\frac{\gamma\,(\gamma+1)}{b_{i}\,l_{i}\,\gamma^{q}}\right)^{\frac{1}{q-1}}.$ (5.5)
Step 3: Liminf–limsup estimates.
The inequalities for $\liminf$ and $\limsup$ in (2.6) follow from the comparison principle applied to the barriers $C_{-}\,v^{-\gamma}$ and $C_{+}\,v^{-\gamma}$ from Lemma 3.2. Since $C_{-}\leq u\,d^{\gamma}\leq C_{+}$ and the constants $C_{\pm}$ can be made arbitrarily close to $C^{*}$ (from above and below), we obtain the precise bounds $\xi_{1}$ and $\xi_{2}$ by using the extremal values $b_{2},l_{2}$ (for $\xi_{1}$, the lower bound) and $b_{1},l_{1}$ (for $\xi_{2}$, the upper bound).
If both limits $b_{0}=\lim_{d\to 0}b(x)\,d^{-\beta}$ and $l_{0}=\lim_{s\to\infty}h(s)\,s^{-q}$ exist, then $b_{1}=b_{2}=b_{0}$ and $l_{1}=l_{2}=l_{0}$, which forces $\xi_{1}=\xi_{2}=\xi_{0}$, establishing (2.8). ∎
Remark 5.1 (The Universal Correction Constant).
The constant $\xi_{0}$ defined in (2.8) has a direct economic interpretation. In the stochastic control framework (Section 6), $\xi_{0}$ determines the intensity of the “boundary penalty” that the optimal strategy must impose to keep the system within $\Omega$. A larger $\xi_{0}$ (corresponding to smaller $b_{0}$ or $l_{0}$, i.e., weaker singular weights or weaker gradient cost) means the controller must exert a stronger corrective effort near the boundary. Conversely, a smaller $\xi_{0}$ (stronger singular weights) means the natural cost structure already provides significant deterrence, reducing the required corrective effort.
6. Stochastic Control and Numerical Verification
In this section, we develop the complete stochastic control framework that gives rise to the mathematical problem studied. We provide all details leading from the controlled diffusion model to the Hamilton–Jacobi–Bellman equation (1.3), and we prove the verification theorem establishing the equivalence between the PDE solution and the value function.
6.1. The Controlled Diffusion Model
Let $(\Omega_{0},\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space supporting an $N$-dimensional standard Brownian motion $\{W_{t}\}_{t\geq 0}$, satisfying the usual conditions (completeness and right-continuity of the filtration).
Consider a system whose state $X_{t}\in\mathbb{R}^{N}$ evolves according to the controlled stochastic differential equation (SDE):
$dX_{t}=\xi_{t}\,dt+\sqrt{2}\,dW_{t},\quad X_{0}=x\in\Omega,$ (6.1)
where $\xi_{t}$ is the control process (representing the corrective action taken by the decision-maker) and the factor $\sqrt{2}$ is chosen so that the generator of the uncontrolled diffusion is exactly the Laplacian $\Delta$ (rather than $\frac{1}{2}\Delta$).
The derivation of the generator proceeds as follows. For a smooth function $\phi\in C^{2}(\Omega)$, Itô’s formula gives:
$d\phi(X_{t})$ $=\nabla\phi(X_{t})\cdot dX_{t}+\frac{1}{2}\,\mathrm{tr}\bigl(D^{2}\phi(X_{t})\,(\sqrt{2}\,I)(\sqrt{2}\,I)^{T}\bigr)\,dt$
$=\bigl[\nabla\phi(X_{t})\cdot\xi_{t}+\Delta\phi(X_{t})\bigr]\,dt+\sqrt{2}\,\nabla\phi(X_{t})\cdot dW_{t}.$ (6.2)
The infinitesimal generator of the uncontrolled process ($\xi\equiv 0$) is therefore $\mathcal{A}\phi=\Delta\phi$, as claimed.
6.2. Admissibility and State Constraints
The fundamental constraint is that the state must remain in $\Omega$ at all times.
Definition 6.1 (Strict admissibility).
A control process $\xi=\{\xi_{t}\}_{t\geq 0}$ is strictly admissible starting from $x$, denoted $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$, if:
  1. (1)
    $\xi$ is progressively measurable with respect to $\{\mathcal{F}_{t}\}$;
  2. (2)
    The SDE (6.1) has a unique strong solution;
  3. (3)
    The process $X_{t}$ remains in $\Omega$ for all time almost surely:
    $\mathbb{P}\bigl(X_{t}\in\Omega\text{ for all }t\geq 0\bigr)=1.$ (6.3)
Remark 6.2.
The set $\mathcal{A}_{\mathrm{sc}}(x)$ is nonempty for each $x\in\Omega$. This is a nontrivial fact that follows from the existence of the blow-up solution and the construction of optimal feedback controls below (Theorem 6.9).
6.3. Legendre Transform and Running Cost
The connection between the PDE (1.3) and stochastic control is mediated by the Legendre–Fenchel transform (convex conjugation).
Definition 6.3 (Convex conjugate).
For $h$ satisfying Assumption 2.4, define the convex conjugate
$h^{*}(s)=\sup_{t\geq 0}\bigl\{s\,t-h(t)\bigr\},\quad s\geq 0.$ (6.4)
By standard convex analysis (see, e.g., Rockafellar [25]), $h^{*}$ has the following properties:
Lemma 6.4 (Properties of the conjugate).
Under Assumption 2.4:
  1. (1)
    $h^{*}$ is strictly convex, $h^{*}(0)=0$, and $h^{*}\in C^{1}(0,\infty)$;
  2. (2)
    $h^{*}(s)\asymp s^{q^{\prime}}$ where $q^{\prime}=q/(q-1)$ is the Hölder conjugate exponent;
  3. (3)
    The supremum in (6.4) is achieved at $t=(h^{\prime})^{-1}(s)$, and $(h^{*})^{\prime}(s)=(h^{\prime})^{-1}(s)$;
  4. (4)
    The Fenchel–Young inequality holds: $s\,t\leq h(t)+h^{*}(s)$ for all $s,t\geq 0$, with equality if and only if $s=h^{\prime}(t)$.
Definition 6.5 (Running cost).
The running cost function is defined as
$L(x,\xi)=b(x)\,h^{*}\left(\frac{|\xi|}{b(x)}\right).$ (6.5)
Lemma 6.6 (Properties of the running cost).
The function $L$ defined in (6.5) satisfies:
  1. (1)
    $L(x,\cdot)$ is strictly convex for each $x\in\Omega$;
  2. (2)
    $L(x,0)=0$;
  3. (3)
    $L(x,\xi)\asymp b(x)^{1-q^{\prime}}\,|\xi|^{q^{\prime}}$ as $|\xi|\to\infty$;
  4. (4)
    Near $\partial\Omega$: $L(x,\xi)\asymp d(x)^{\beta(1-q^{\prime})}\,|\xi|^{q^{\prime}}$, which blows up as $d(x)\to 0$ for fixed $|\xi|>0$ provided $\beta>0$.
Proof.
Properties (1)–(3) follow directly from the properties of $h^{*}$ (Lemma 6.4) and the positive homogeneity of the scaling $|\xi|/b(x)$. For property (4), we use $b(x)\asymp d(x)^{\beta}$ from Assumption 2.6 (W2) and compute:
$L(x,\xi)\asymp b(x)\left(\frac{|\xi|}{b(x)}\right)^{q^{\prime}}=b(x)^{1-q^{\prime}}\,|\xi|^{q^{\prime}}\asymp d(x)^{\beta(1-q^{\prime})}\,|\xi|^{q^{\prime}}.$
Since $q^{\prime}>1$, we have $1-q^{\prime}<0$, so $\beta(1-q^{\prime})<0$ when $\beta>0$, ensuring $L(x,\xi)\to\infty$ as $d(x)\to 0$.
The crucial duality identity is: for each $x\in\Omega$,
$b(x)\,h(|p|)=\sup_{\xi\in\mathbb{R}^{N}}\bigl\{-\xi\cdot p-L(x,\xi)\bigr\}.$ (6.6)
This follows from the definition of $h^{*}$ and the substitution $\xi=-b(x)\,h^{\prime}(|p|)\,p/|p|$. ∎
6.4. Derivation of the Hamilton–Jacobi–Bellman Equation
We now derive the HJB equation (1.3) from the stochastic control problem via the dynamic programming principle. This derivation makes explicit how each term in the PDE arises from the control problem.
Definition 6.7 (Value function).
For $x\in\Omega$, define the value function
$V(x)=\inf_{\xi\in\mathcal{A}_{\mathrm{sc}}(x)}\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a(X_{s})\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$ (6.7)
Remark 6.8 (Well-posedness of the cost functional).
The integral in (6.7) is well-defined because:
  1. (1)
    The discount factor $e^{-\int_{0}^{t}a(X_{s})\,ds}$ ensures integrability for large $t$ since $a(x)>0$ everywhere in $\Omega$ (Assumption 2.6).
  2. (2)
    For strictly admissible controls, $X_{t}$ stays in $\Omega$, so $L(X_{t},\xi_{t})$ remains finite for each $t$.
  3. (3)
    The blow-up of $L$ near $\partial\Omega$ (Lemma 6.6, property (4)) creates an infinite penalty for trajectories approaching the boundary, effectively enforcing the state constraint.
Derivation via the Dynamic Programming Principle. Assume $V$ is smooth. For a small time increment $\Delta t>0$, the dynamic programming principle gives:
$V(x)=\inf_{\xi}\,\mathbb{E}\left[\int_{0}^{\Delta t}e^{-\int_{0}^{t}a(X_{s})\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt+e^{-\int_{0}^{\Delta t}a(X_{s})\,ds}\,V(X_{\Delta t})\right].$ (6.8)
Expanding $V(X_{\Delta t})$ via Itô’s formula (6.2) and taking $\Delta t\to 0$, we obtain:
$0=\inf_{\xi\in\mathbb{R}^{N}}\left\{\Delta V(x)+\nabla V(x)\cdot\xi-a(x)\,V(x)+L(x,\xi)+f(x)\right\}.$ (6.9)
The infimum over $\xi$ can be computed explicitly using (6.6). Using the identity $\inf_{\xi}\{\nabla V\cdot\xi+L(x,\xi)\}=-\sup_{\xi}\{-\nabla V\cdot\xi-L(x,\xi)\}=-b(x)\,h(|\nabla V|)$, we obtain:
$0=\Delta V-b(x)\,h(|\nabla V|)-a(x)\,V+f(x),$ (6.10)
which is precisely equation (1.3) with $u=V$.
This derivation shows explicitly how each term in (1.3) arises:
  • $\Delta V$: the generator of the Brownian noise $\sqrt{2}\,dW_{t}$;
  • $b(x)\,h(|\nabla V|)$: the optimal cost of corrective action, via Legendre duality;
  • a(x)Va(x)\,V: the state-dependent discounting;
  • $f(x)$: the external source/cost.
6.5. Verification Theorem
Theorem 6.9 (Verification Theorem — Stochastic Representation).
Under Assumptions 2.12.6, the unique solution $u\in C^{2}(\Omega)$ from Theorem 2.9 coincides with the value function:
$u(x)=V(x)\quad\text{for all }x\in\Omega.$ (6.11)
Moreover, the optimal feedback control is given by
$\xi^{*}(x)=-b(x)\,h^{\prime}(|\nabla u(x)|)\,\frac{\nabla u(x)}{|\nabla u(x)|}=-b(x)\,(h^{*})^{\prime}(|\nabla u(x)|)^{-1}\,\frac{\nabla u(x)}{|\nabla u(x)|}.$ (6.12)
The identity $(h^{\prime})^{-1}=(h^{*})^{\prime}$ is a standard consequence of the Legendre transform: the supremum in $h^{*}(s)=\sup_{t}\{st-h(t)\}$ is achieved at $s=h^{\prime}(t)$, so $t=(h^{\prime})^{-1}(s)$. Differentiating the identity $h^{*}(s)=s\,(h^{\prime})^{-1}(s)-h((h^{\prime})^{-1}(s))$ with respect to $s$ yields $(h^{*})^{\prime}(s)=(h^{\prime})^{-1}(s)$.
Proof.
The proof consists of two parts: establishing $u\leq V$ (sub-optimality) and $u\geq V$ (optimality).
Part 1: Sub-optimality ($u\leq V$).
Let $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ be any strictly admissible control, and let $X_{t}$ be the corresponding trajectory satisfying (6.1). For $n\in\mathbb{N}$, define the stopping times
$\tau_{n}=\inf\{t>0:d(X_{t})\leq 1/n\}\wedge n.$ (6.13)
Since $\xi$ is strictly admissible, $X_{t}$ remains in $\Omega$ almost surely for all $t\geq 0$. For any $T>0$, the path $t\mapsto X_{t}$ is continuous and the image $X([0,T])$ is a compact subset of $\Omega$, hence $\inf_{t\in[0,T]}d(X_{t})>0$ almost surely. This ensures that for almost all $\omega$, there exists $n_{0}(\omega)$ such that $\tau_{n}(\omega)=n$ for all $n\geq n_{0}(\omega)$, giving $\tau_{n}\to\infty$ as $n\to\infty$ almost surely.
Apply Itô’s formula to $e^{-\int_{0}^{t}a(X_{s})\,ds}\,u(X_{t})$ on $0,\tau_{n}]$. By (6.2):
$e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})$
$=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl[-a(X_{t})\,u(X_{t})+\Delta u(X_{t})+\nabla u(X_{t})\cdot\xi_{t}\bigr]\,dt$
$\quad+\sqrt{2}\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\nabla u(X_{t})\cdot dW_{t}.$ (6.14)
Using the PDE $\Delta u+b\,h(|\nabla u|)+a\,u=f$, we substitute $\Delta u=b\,h(|\nabla u|)+a\,u-f$. Thus (6.14) becomes:
$e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})$ $=u(x)+\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl[b\,h(|\nabla u|)-f+\nabla u\cdot\xi_{t}\bigr]\,dt$
$\quad+\text{martingale}.$ (6.15)
By the Fenchel–Young inequality (Lemma 6.4, property (4)):
$b(x)\,h(|p|)+p\cdot\xi\geq-L(x,\xi)\quad\text{for all }p,\xi\in\mathbb{R}^{N}.$ (6.16)
Applying (6.16) with $p=\nabla u(X_{t})$:
$b\,h(|\nabla u|)+\nabla u\cdot\xi_{t}\geq-L(X_{t},\xi_{t}).$
Taking expectations in (6.15) (the Itô integral is a martingale with zero expectation, since $u$ and $\nabla u$ are bounded on the compact set $\{x:d(x)\geq 1/n\}$):
$\mathbb{E}\bigl[e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})\bigr]\geq u(x)-\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$ (6.17)
Rearranging:
$u(x)\leq\mathbb{E}\left[\int_{0}^{\tau_{n}}e^{-\int_{0}^{t}a\,ds}\,\bigl(L+f\bigr)\,dt\right]+\mathbb{E}\bigl[e^{-\int_{0}^{\tau_{n}}a\,ds}\,u(X_{\tau_{n}})\bigr].$ (6.18)
As $n\to\infty$: the first term converges to the full integral by monotone convergence. For the second term, since $\xi$ is strictly admissible, $X_{\tau_{n}}$ stays in $\Omega$, so $u(X_{\tau_{n}})$ is finite. Moreover, $e^{-\int_{0}^{\tau_{n}}a\,ds}\to 0$ as $\tau_{n}\to\infty$ (since $a>0$), so this term vanishes.
Taking $n\to\infty$ in (6.18):
$u(x)\leq\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t},\xi_{t})+f(X_{t})\bigr)\,dt\right].$
Since $\xi\in\mathcal{A}_{\mathrm{sc}}(x)$ was arbitrary, $u(x)\leq V(x)$.
Part 2: Optimality ($u\geq V$).
Consider the feedback control defined by (6.12). At this choice, the Fenchel–Young inequality (6.16) becomes an equality:
$b\,h(|\nabla u|)+\nabla u\cdot\xi^{*}=-L(x,\xi^{*}).$ (6.19)
This is the first-order optimality condition for the Legendre transform: the supremum in $h^{*}$ is achieved.
Let $X_{t}^{*}$ be the trajectory under the feedback control $\xi_{t}^{*}=\xi^{*}(X_{t}^{*})$. The blow-up behavior of $u$ near $\partial\Omega$ implies $|\xi^{*}(x)|\to\infty$ as $x\to\partial\Omega$. Specifically, since $\nabla u$ blows up as $d(x)^{-\gamma-1}$, the drift $\xi^{*}$ points strongly into the interior of $\Omega$ near the boundary.
By Lemma 1.1 of Lasry–Lions [17], this singular drift prevents the process from reaching $\partial\Omega$:
$\mathbb{P}\bigl(X_{t}^{*}\in\Omega\text{ for all }t\geq 0\bigr)=1.$
Thus $\xi^{*}\in\mathcal{A}_{\mathrm{sc}}(x)$.
Repeating the Itô calculation with equality in (6.16) (i.e., using (6.19)):
$u(x)=\mathbb{E}\left[\int_{0}^{\infty}e^{-\int_{0}^{t}a\,ds}\,\bigl(L(X_{t}^{*},\xi_{t}^{*})+f(X_{t}^{*})\bigr)\,dt\right]\geq V(x).$
Combining Parts 1 and 2: $u(x)=V(x)$. ∎
Remark 6.10 (Connection to Lasry–Lions Theorem 1.1).
Theorem 6.9 generalizes Theorem 1.1 in [17] to equations with state-dependent weight functions $a(x)$ and $b(x)$. The original Lasry–Lions result treats the case $a=\lambda\in(0,\infty)$ (constant discount rate), $b=1$ (no spatial weight), and $h(|\nabla u|)=|\nabla u|^{p}$ for $p>1$. Our formulation incorporates:
  1. (1)
    A state-dependent discount rate $a(x)$ in the cost functional;
  2. (2)
    A state-dependent control cost coefficient $b(x)$;
  3. (3)
    A general strictly convex Hamiltonian $h$ with power-type growth;
  4. (4)
    A running reward/cost $f(x)$.
The key insight of Lasry–Lions — that the blow-up of the value function enforces state constraints through the singularity of the optimal drift — remains valid in this more general setting.
Remark 6.11 (Geometric intuition of the optimal drift).
Following the seminal work of Lasry and Lions [17], we provide a rigorous geometric explanation for why the optimal drift $\xi^{*}(x)$ is oriented towards the interior of $\Omega$.
As $x$ approaches $\partial\Omega$, the blow-up condition $u(x)\to+\infty$ implies that $\nabla u(x)$ points in the direction of steepest increase, which is the outward normal direction. Near $\partial\Omega$, the approximation $u(x)\approx C\,d(x)^{-\gamma}$ yields:
$\nabla u(x)\approx-C\,\gamma\,d(x)^{-\gamma-1}\,\nabla d(x),$
where $\nabla d(x)$ points into the interior; thus $\nabla d(x)$ points towards the boundary, and $\nabla u$ points towards the boundary (outward).
The optimal feedback control (6.12) has the form $\xi^{*}(x)\propto-\nabla u(x)$, which therefore points into the interior. The magnitude $|\xi^{*}(x)|\asymp d(x)^{-(\gamma+1)(q-1)}$ blows up as $d(x)\to 0$, creating a singular inward drift that overwhelms the Brownian noise and prevents the process from reaching $\partial\Omega$.
Remark 6.12 (Alternative proof of convexity via stochastic control).
A purely control-theoretic justification for the convexity of $u$ follows from the stochastic representation $u=V$ and the results of Alvarez [1]. Since the controlled diffusion in (6.1) is linear in $(x,\xi)$ and the running cost $L(x,\xi)$ is jointly convex in $(x,\xi)$ by our structural assumptions on the perspective function of $h^{*}$, the value function $V(x)$ is the infimum of a jointly convex functional over a convex set of control–trajectory pairs. In an alternative framework, under the convexity of $\Omega$ and the assumptions on $a$ and $b$, together with $f$ convex (so that the total running cost $L(x,\xi)+f(x)$ remains jointly convex in $(x,\xi)$), the results of [1] imply the convexity of $u(x)$. This provides a probabilistic verification that is complementary to the microscopic convexity principle used in Section 4.
7. Applications in Operational Research and Management Science
In this section, we develop a comprehensive framework for applying the mathematical results of Sections 46 to problems arising in Operational Research (OR) and Management Science. We present three detailed case studies and explore interdisciplinary connections with physics.
7.1. General Framework for Economic Applications
The stochastic control problem (6.1)–(6.7) provides a natural framework for modeling systems that must operate within bounded feasible regions. The key elements of this framework are:
Math Symbol Economic Meaning Physical Analogue
$X_{t}\in\Omega$ System state (inventory, portfolio) Particle position
$\xi_{t}$ Control action (reorder, rebalance) External force
$\sqrt{2}\,dW_{t}$ Demand/market noise Thermal fluctuation
$a(x)$ State-dependent discount rate Damping coefficient
$b(x)$ Adjustment stiffness coefficient Viscosity
$h(|\nabla u|)$ Convex adjustment cost Kinetic energy
$u(x)=V(x)$ Optimal cost-to-go Free energy
$\partial\Omega$ Constraint boundary Potential wall
$u\to\infty$ at $\partial\Omega$ Infinite penalty Infinite barrier
$\gamma$ Blow-up rate Critical exponent
$\xi_{0}$ Universal correction constant Scaling amplitude
The table above reveals a deep structural analogy between economic optimization, mathematical analysis, and statistical physics. In all three domains, the solution to the boundary-value problem encodes the optimal response of a system to stochastic perturbations under hard constraints.
7.2. Case Study 1: Multi-Product Inventory Management
Problem Description. Consider a warehouse managing $N$ distinct products. Let $X_{t}=(X_{t}^{1},\ldots,X_{t}^{N})$ denote the vector of inventory levels at time $t$. Each product must satisfy:
$X_{t}^{i}\in(L_{i},U_{i}),\quad i=1,\ldots,N,$ (7.1)
where $L_{i}$ is the minimum stock level (below which stockouts occur) and $U_{i}$ is the maximum storage capacity. The feasible region is the hyperrectangle $\Omega=\prod_{i=1}^{N}(L_{i},U_{i})$, which is a convex domain satisfying Assumption 2.1.
Model Formulation. Each product’s inventory evolves as:
$dX_{t}^{i}=\xi_{t}^{i}\,dt+\sigma_{i}\,dW_{t}^{i},$ (7.2)
where $\xi_{t}^{i}$ is the production/procurement rate for product $i$ (the control), $\sigma_{i}$ captures demand uncertainty, and $W_{t}^{i}$ are independent Brownian motions representing stochastic demand fluctuations.
Cost Structure. The adjustment cost has the structure:
$L(x,\xi)=b(x)\,h^{*}\left(\frac{|\xi|}{b(x)}\right),$ (7.3)
where:
  • $b(x)=\prod_{i=1}^{N}(x^{i}-L_{i})^{\beta_{i}/N}\cdot(U_{i}-x^{i})^{\beta_{i}/N}$ models the increasing difficulty of adjustments near capacity limits. As inventory approaches either bound, the physical constraints of the warehouse (loading dock congestion, emergency procurement lead times) increase the cost per unit adjustment.
  • $a(x)=\sum_{i=1}^{N}c_{i}\,(x^{i}-L_{i})^{\alpha_{i}}$ is the holding cost rate, which increases as inventory approaches the lower bound (opportunity cost of potential stockouts).
  • $f(x)$ represents the baseline operating cost, assumed constant or concave (reflecting economies of scale).
Application of Main Results. By Theorem 2.9, the optimal cost-to-go function $V(x)$ exists, is unique, and blows up at the boundary of the feasible region. The blow-up rate $\gamma=(\beta-q+2)/(q-1)$ quantifies how rapidly the cost of maintaining feasibility increases near the constraints.
By Theorem 2.11, the value function $V(x)$ is strictly convex, which guarantees that the optimal reorder policy $\xi^{*}(x)$ (given by (6.12)) is uniquely determined for each inventory state. This is essential for implementation: the warehouse manager can compute the optimal action unambiguously.
The Universal Correction Constant $\xi_{0}$ from Theorem 2.10 determines the “safety stock” level: it quantifies the intensity of the preventive corrective action that the optimal policy imposes near the constraints.
Numerical implications. For the $N=1$ case (single product), the problem reduces to a one-dimensional domain $\Omega=(L,U)$. For $N=2$ (two products), we obtain a two-dimensional domain, and the numerical simulations in Section 9 directly illustrate the structure of the optimal policy.
7.3. Case Study 2: Portfolio Risk Management under VaR Constraints
Problem Description. Consider a portfolio manager allocating wealth across $N$ asset classes. Let $X_{t}=(X_{t}^{1},\ldots,X_{t}^{N})$ denote the vector of portfolio weights. The weights must satisfy:
$X_{t}^{i}\geq 0\quad\text{(no short selling)},\quad\sum_{i=1}^{N}X_{t}^{i}\leq 1\quad\text{(budget constraint)}.$ (7.4)
The feasible region is the standard simplex $\Omega=\{x\in\mathbb{R}^{N}:x^{i}>0,\,\sum x^{i}<1\}$, which is a bounded convex domain.
Model Formulation. The portfolio weights evolve under a rebalancing strategy:
$dX_{t}^{i}=\xi_{t}^{i}\,dt+\sum_{j=1}^{N}\sigma_{ij}(X_{t})\,dW_{t}^{j},$ (7.5)
where $\xi_{t}^{i}$ represents the rebalancing rate and $\sigma_{ij}$ captures the diffusion driven by asset returns.
Connection to the Framework. The HJB equation (1.3) governs the value function $V(x)$, which represents the minimum expected rebalancing cost subject to the constraint that the portfolio weights remain in the simplex. The singular weight $b(x)$ captures the increasing transaction costs near the constraint boundaries (e.g., market impact costs for large trades when a position is near zero or the full budget).
By Theorem 2.10, the optimal rebalancing intensity near the constraint boundaries is governed by the Universal Correction Constant $\xi_{0}$. This provides a quantitative formula for the “buffer zone” that risk-averse managers should maintain.
Risk management implications. The convexity of $V(x)$ (Theorem 2.11) ensures that the optimal rebalancing policy is stabilizing: any perturbation of the portfolio weights triggers a unique corrective action that drives the system back toward the interior of the feasible region. This is the mathematical formulation of the “mean-reverting” behavior that is desirable in portfolio management.
The blow-up rate $\gamma$ can be interpreted as a measure of the “regulatory sensitivity” of the portfolio: higher $\gamma$ means that the cost of approaching constraint violations increases more rapidly, providing stronger deterrence.
7.4. Case Study 3: Supply Chain Network Optimization
Problem Description. Consider a supply chain with $N$ nodes (warehouses, factories, distribution centers). The state $X_{t}=(X_{t}^{1},\ldots,X_{t}^{N})$ represents the inventory or throughput at each node. The system must satisfy capacity constraints at each node and flow conservation requirements between nodes.
Interaction with Physics: Fluid Dynamics Analogy. The supply chain optimization problem has a direct analogy with fluid dynamics. The material flow in the supply chain behaves like a fluid in a network of pipes:
  • Nodes correspond to reservoirs with finite capacity;
  • Edges correspond to pipes with finite throughput;
  • Demand noise corresponds to turbulent fluctuations;
  • Optimal control corresponds to pressure regulation.
The HJB equation (1.3) in this context is analogous to the pressure equation in incompressible fluid dynamics. The boundary blow-up condition $u\to\infty$ at $\partial\Omega$ corresponds to the infinite pressure required to prevent the fluid from escaping the network — a direct analogue of the “no-flow” boundary condition in fluid mechanics.
The boundary layer connection. In fluid dynamics, the behavior of a viscous fluid near a solid boundary is described by the boundary layer theory of Prandtl. The velocity profile transitions from zero at the wall (no-slip condition) to the free-stream velocity over a thin layer of width $\delta\sim\text{Re}^{-1/2}$, where Re is the Reynolds number.
In our mathematical framework, an analogous “boundary layer” exists. The solution $u(x)$ transitions from moderate values in the interior to infinite values at $\partial\Omega$ over a layer whose effective width is determined by the blow-up rate $\gamma$:
$\delta_{\text{eff}}\sim\left(\frac{C_{\text{interior}}}{C_{\text{boundary}}}\right)^{1/\gamma}.$ (7.6)
This “economic boundary layer” determines the region near the constraints where the optimal policy transitions from steady-state operation to emergency corrective action.
7.5. Interactions with Physics: Diffusion and Statistical Mechanics
The mathematical structure of our problem has deep connections with several areas of physics.
Connection 1: Brownian motion and diffusion. The controlled SDE (6.1) describes a particle undergoing Brownian motion with an external drift. The value function $V(x)=u(x)$ plays the role of the free energy of the system:
$u(x)=-k_{B}T\,\ln Z(x),$ (7.7)
where $Z(x)$ is the partition function and $k_{B}T$ is the thermal energy. The boundary blow-up $u\to\infty$ corresponds to $Z\to 0$: near the boundary, the number of accessible microstates vanishes, creating an “entropic barrier” that confines the system.
Connection 2: Burgers equation and nonlinear waves. Through the Hopf–Cole transformation $u=-2\nu\,\ln\psi$, the viscous Burgers equation can be related to the heat equation. Similarly, our equation (1.3) with $h(s)=s^{2}$ can be transformed into a linear equation via an exponential substitution. The blow-up at the boundary corresponds to the formation of a shock wave in the Burgers equation — a discontinuity where the gradient becomes infinite. In the economic interpretation, this corresponds to the sudden transition from normal operations to crisis mode.
Connection 3: Optimal transport. The Monge-Ampère equation studied by Zhang [27] arises naturally in the theory of optimal transport. The singular weights in our equation play a role analogous to the cost function in the transport problem: they determine the “difficulty” of moving mass from one location to another. Our extension to the Laplacian setting (from the Monge-Ampère setting) broadens the applicability to problems with stochastic perturbations.
7.6. Practical Implementation Guidelines for Managers
Based on the theoretical results, we provide the following guidelines for practitioners in Operational Research:
  1. (1)
    Identify the feasible region $\Omega$. Define the hard constraints on the system state. Verify that the feasible region is convex (a natural property for most operational constraints).
  2. (2)
    Estimate the singular exponents. Determine $\beta$ (the rate at which adjustment costs increase near constraints) and $q$ (the power-law exponent of the adjustment cost function) from empirical data.
  3. (3)
    Compute the blow-up rate. Use $\gamma=(\beta-q+2)/(q-1)$ to determine which asymptotic regime applies. If $\gamma>0$ (gradient-dominant), the theory provides explicit asymptotic formulas.
  4. (4)
    Calibrate the Universal Correction Constant. Use $\xi_{0}=(\gamma(\gamma+1)/(b_{0}\,l_{0}\,\gamma^{q}))^{1/(q-1)}$ to determine the optimal intensity of corrective action near constraints.
  5. (5)
    Implement the feedback control. The optimal policy $\xi^{*}(x)=-b(x)\,h^{\prime}(|\nabla u(x)|)\,\nabla u(x)/|\nabla u(x)|$ can be computed numerically using the monotone iterative scheme of Section 8.
8. Monotone Iterative Construction
For numerical implementation and as an alternative existence proof, we develop a monotone iterative scheme that converges to the unique solution.
8.1. The Iterative Scheme
Let $\Lambda:\Omega\to(0,\infty)$ be a strictly positive weight function satisfying
$\Lambda(x)\geq\max\left\{a(x),\,b(x)\,q\,l_{2}\,(C_{+}\,\gamma)^{q-1}\,v(x)^{-(\gamma+1)(q-1)+\beta}\right\}.$ (8.1)
This choice ensures monotonicity of the iteration by making the right-hand side of the linearized equation an increasing function of the iterates.
Definition 8.1 (Monotone iteration).
Define the sequence $\{u^{(k)}\}_{k=0}^{\infty}$ as follows:
  1. (1)
    Initialization: $u^{(0)}(x)=C_{-}\,v(x)^{-\gamma}$ (the subsolution from Lemma 3.2).
  2. (2)
    Iteration: For $k\geq 0$, let $u^{(k+1)}$ be the unique solution to
    $\left\{\begin{array}[]{ll}-\Delta u^{(k+1)}+\Lambda(x)\,u^{(k+1)}=\mathcal{F}(x,u^{(k)},\nabla u^{(k)})&\text{in }\Omega_{\delta},\\ u^{(k+1)}=C_{+}\,v^{-\gamma}&\text{on }\partial\Omega_{\delta},\end{array}\right$. (8.2)
    where $\Omega_{\delta}=\{x\in\Omega:d(x)>\delta\}$ and
    $\mathcal{F}(x,u,p)=\Lambda(x)\,u-b(x)\,h(|p|)-a(x)\,u+f(x).$ (8.3)
Theorem 8.2 (Convergence of Monotone Iteration).
Under the assumptions of Theorem 2.9, the sequence $\{u^{(k)}\}$ satisfies:
  1. (1)
    Monotonicity: $u^{(0)}\leq u^{(1)}\leq\cdots\leq u^{(k)}\leq\cdots\leq\overline{u}$;
  2. (2)
    Convergence: $u^{(k)}\to u$ uniformly on compact subsets of $\Omega$;
  3. (3)
    Rate: For fixed compact $K\Subset\Omega$, there exist $C,\rho>0$ with $\rho<1$ such that
    $\|u^{(k)}-u\|_{L^{\infty}(K)}\leq C\,\rho^{k}.$
Proof.
Step 1: Monotonicity.
We show $u^{(k)}\leq u^{(k+1)}$ by induction. For $k=0$: the function $w=u^{(1)}-u^{(0)}$ satisfies
$-\Delta w+\Lambda\,w$ $=\mathcal{F}(x,u^{(0)},\nabla u^{(0)})-(-\Delta u^{(0)}+\Lambda\,u^{(0)})$
$=-\mathcal{L}u^{(0)}\geq 0,$
since $u^{(0)}$ is a subsolution (i.e., $\mathcal{L}u^{(0)}\leq 0$). By the maximum principle for the operator $\Delta+\Lambda$ (with $\Lambda>0$), $w\geq 0$, i.e., $u^{(1)}\geq u^{(0)}$.
For the inductive step, suppose $u^{(k)}\geq u^{(k-1)}$. The right-hand side $\mathcal{F}(x,u,p)$ is monotone increasing in $u$ since $\partial\mathcal{F}/\partial u=\Lambda(x)-a(x)\geq 0$ by (8.1). Combined with careful handling of the gradient dependence through the choice of $\Lambda$, the iteration operator is order-preserving.
Step 2: Upper bound.
Similarly, $u^{(k)}\leq\overline{u}$ for all $k$ follows from the supersolution property of $\overline{u}$ and the same monotonicity argument.
Step 3: Convergence.
The sequence $\{u^{(k)}\}$ is monotone increasing and bounded above by $\overline{u}$. By Dini’s theorem, convergence is uniform on compact sets. Let $u_{\infty}=\lim_{k\to\infty}u^{(k)}$.
By interior Schauder estimates, $\{u^{(k)}\}$ is bounded in $C^{2,\alpha}(K)$ for any compact $K\Subset\Omega$. By the Arzelà–Ascoli theorem, a subsequence converges in $C^{2}(K)$. Since the full sequence is monotone, the entire sequence converges in $C^{2}_{\mathrm{loc}}(\Omega)$.
Passing to the limit in (8.2), we find $\mathcal{L}u_{\infty}=0$ in $\Omega_{\delta}$. Since this holds for all $\delta>0$, $u_{\infty}$ solves $\mathcal{L}u_{\infty}=0$ in $\Omega$. By uniqueness (Theorem 2.9), $u_{\infty}=u$.
Step 4: Convergence rate.
The linear convergence rate follows from the contraction mapping principle in appropriate weighted spaces; see [2] for general monotone iteration theory. ∎
8.2. Discretization and Implementation
For numerical implementation, the iteration (8.2) is discretized on a mesh. Key considerations include:
  • Variational formulation: The linear problem (8.2) admits the weak form: find $u^{(k+1)}\in H^{1}(\Omega_{\delta})$ such that
    $\int_{\Omega_{\delta}}\left(\nabla u^{(k+1)}\cdot\nabla\phi+\Lambda\,u^{(k+1)}\,\phi\right)dx=\int_{\Omega_{\delta}}\mathcal{F}(u^{(k)})\,\phi\,dx$ (8.4)
    for all test functions $\phi\in H_{0}^{1}(\Omega_{\delta})$.
  • Regularized boundary: The blow-up boundary condition is enforced on the truncated domain $\partial\Omega_{\delta}$ by applying the Dirichlet condition $u=C_{+}\,\delta^{-\gamma}$ for small $\delta>0$.
  • Relaxation: For improved stability, use the damped update $u^{(k+1)}_{\text{new}}=\omega\,u^{(k+1)}+(1-\omega)\,u^{(k)}$ with $\omega\in(0,1]$.
9. Numerical Simulations for $N=1$ and $N=2$
In this section, we present comprehensive numerical simulations that validate the theoretical predictions of Sections 45 and illustrate the economic applications of Section 7. We consider both the one-dimensional case ($N=1$) and the two-dimensional case ($N=2$), using the monotone iterative scheme of Section 8. The complete Python codes are provided in Appendices A and B.
9.1. One-Dimensional Simulations ($N=1$)
9.1.1. Setup
We use the domain $\Omega=(-1,1)$ with the defining function $v(x)=1-x^{2}$, which is strictly concave on (1,1)(-1,1). The parameters are:
  • Gradient exponent: $q=1.6$ (gradient-dominant regime, since $q=1.6<2.5=\beta+2$);
  • Weight exponent: $\beta=0.5$;
  • Reaction exponent: $\alpha=-0.2$;
  • Source term: $f(x)\equiv 1.0$;
  • Grid: $N=400$ points on [0.95,0.95][-0.95,0.95] (truncation $\delta=0.05$).
The theoretical blow-up exponent is:
$\gamma=\frac{\beta-q+2}{q-1}=\frac{0.5-1.6+2}{1.6-1}=\frac{0.9}{0.6}=1.5.$ (9.1)
9.1.2. Verification of Blow-up Rates (Figure 1)
Figure 1 compares the numerical solutions in the three asymptotic regimes identified in Lemma 3.2:
  • Case 1 (Gradient-dominant): $q=1.6$, $\beta=0.5$, giving $\gamma=1.5$. The solution exhibits power-law blow-up $u(x)\sim C\,d(x)^{-1.5}$ near $x=\pm 1$.
  • Case 2 (High-order): $q=3.0$, $\beta=0.5$. Here $q>\beta+2=2.5$, so the gradient term dominates at all orders. The solution still blows up but with a different (steeper) profile.
  • Case 3 (Logarithmic): $q=2.5$, $\beta=0.5$. Since $q=\beta+2$, this is the critical case with logarithmic blow-up $u(x)\sim C\,\ln(1/d(x))$.
Figure 1. Comparison of the three boundary blow-up regimes for $\beta=0.5$: gradient-dominant ($q=1.6$, $\gamma=1.5$), high-order ($q=3.0$), and critical logarithmic ($q=2.5$). The solution profiles confirm distinct asymptotic behaviors near $x=\pm 1$. Economic interpretation: In the inventory management context (Section 7.2), the three profiles represent different cost structures for emergency corrective action. The gradient-dominant case (mildest blow-up) corresponds to moderate adjustment costs; the high-order case (steepest blow-up) corresponds to extremely stiff adjustment mechanisms; the logarithmic case is the critical transition between the two regimes.
9.1.3. Scale Analysis (Figure 2)
Figure 2 presents rigorous scale analysis confirming the theoretical predictions:
  • Log-log plot for Case 1: Plotting $\log u$ versus $\log d(x)$ yields a straight line with slope $\gamma$. The linear regression gives a measured slope that matches the theoretical value $\gamma=-1.5$ with high precision ($R^{2}>0.99$). This confirms the power-law blow-up $u\sim C\,d^{-\gamma}$.
  • Semi-log plot for Case 3: Plotting $u$ versus $\log(1/d)$ yields a straight line, confirming the logarithmic blow-up $u\sim C\,\ln(1/d)$.
  • Gradient magnitude comparison: The gradient $|\nabla u|$ diverges as $d^{-\gamma-1}$ in Case 1 and as $d^{-1}$ in Case 3. The different rates confirm the distinct asymptotic regimes.
Figure 2. Scale analysis confirming theoretical blow-up rates for $\beta=0.5$. Top-left: log-log plot for Case 1 showing slope $\gamma=-1.5$. Top-right: solution profile for Case 1. Bottom-left: semi-log plot for Case 3 showing linear growth in $\log(1/d)$. Bottom-right: gradient magnitude comparison between Cases 1 and 3 on logarithmic scale.
9.1.4. Theoretical Consistency Verification (Figure 3)
Figure 3 provides comprehensive verification of all theoretical predictions for the gradient-dominant case ($q=1.6$, $\beta=0.5$):
  1. (a)
    Solution between barriers: The numerical solution $u(x)$ lies between the subsolution $C_{-}\,v^{-\gamma}$ and the supersolution $C_{+}\,v^{-\gamma}$, confirming Theorem 2.9. The plot uses logarithmic scale to visualize the blow-up behavior.
  2. (b)
    Strict convexity: The second derivative $u^{\prime\prime}(x)>0$ throughout (1,1)(-1,1), confirming Theorem 2.11. The convexity is strongest near the boundaries, consistent with the Hessian formula (3.6).
  3. (c)
    Optimal drift structure: The optimal drift $\xi^{*}(x)=-h^{\prime}(|u^{\prime}|)\,\mathrm{sgn}(u^{\prime})\,b(x)$ exhibits the expected antisymmetric structure: it points to the right ($\xi^{*}>0$) for $x<0$ and to the left ($\xi^{*}<0$) for $x>0$, always pushing the state toward the center of $\Omega$. The magnitude becomes singular at the boundaries, confirming Remark 6.11. In the inventory management context, this represents the “emergency reorder” policy that activates as stock levels approach critical bounds.
  4. (d)
    HJB residual: The residual $|\mathcal{L}u|=|-\Delta u+b\,h(|\nabla u|)+a\,u-f|$ is below the convergence tolerance $10^{-6}$ throughout the domain (except near the boundary truncation), confirming that the numerical solution accurately satisfies the PDE.
Figure 3. Theoretical consistency verification for $q=1.6$, $\beta=0.5$: (a) numerical solution bounded by barriers (log scale), (b) strict convexity via $u^{\prime\prime}>0$, (c) singular optimal drift structure, (d) HJB equation residual verification.
9.2. Two-Dimensional Simulations ($N=2$)
9.2.1. Setup
For the two-dimensional case, we use the unit disk $\Omega=\{(x_{1},x_{2})\in\mathbb{R}^{2}:x_{1}^{2}+x_{2}^{2}<1\}$ with the defining function $v(x)=1-|x|^{2}=1-x_{1}^{2}-x_{2}^{2}$, which is strictly concave on $\Omega$ and satisfies $v=0$ on $\partial\Omega$.
The parameters are the same as in the 1D case:
  • $q=1.6$, $\beta=0.5$, $\alpha=-0.2$, $f\equiv 1.0$;
  • $\gamma=1.5$;
  • Grid: $100\times 100$ Cartesian grid on $1,1]^{2}$, masked to the disk interior with truncation $\delta=0.05$.
9.2.2. Economic Interpretation: Two-Product Inventory Problem
In the context of Case Study 1 (Section 7.2), the 2D simulation represents a warehouse managing two products simultaneously. The axes $x_{1}$ and $x_{2}$ represent the normalized inventory levels of products 1 and 2, respectively. The disk constraint $x_{1}^{2}+x_{2}^{2}<1$ can be interpreted as a combined capacity constraint: the total “resource usage” (measured by the Euclidean norm) must not exceed the warehouse capacity.
The value function $V(x_{1},x_{2})$ represents the minimum expected cost of managing both products simultaneously, accounting for demand noise, reorder costs, and the constraint that the combined inventory must remain within the feasible region.
9.2.3. Numerical Method for $N=2$
The 2D solver uses the same monotone iterative scheme (Definition 8.1), discretized on a Cartesian grid with the standard 5-point Laplacian stencil:
$\Delta u\approx\frac{u_{i+1,j}+u_{i-1,j}+u_{i,j+1}+u_{i,j-1}-4\,u_{i,j}}{(\Delta x)^{2}}.$ (9.2)
The gradient magnitude is approximated by central differences:
$|\nabla u|_{i,j}\approx\sqrt{\left(\frac{u_{i+1,j}-u_{i-1,j}}{2\,\Delta x}\right)^{2}+\left(\frac{u_{i,j+1}-u_{i,j-1}}{2\,\Delta x}\right)^{2}}.$ (9.3)
Only grid points satisfying $x_{1}^{2}+x_{2}^{2}<(1-\delta)^{2}$ are included in the computation; boundary points are set to the theoretical blow-up value $C^{*}\,v_{\text{bnd}}^{-\gamma}$.
9.2.4. Results for $N=2$ (Figure 4)
Figure 4 presents the two-dimensional results:
  1. (a)
    Surface plot of $u(x_{1},x_{2})$: The value function exhibits rotational symmetry (as expected for the unit disk with rotationally symmetric weights) and blows up uniformly as $|(x_{1},x_{2})|\to 1$. The minimum is achieved at the origin, which represents the optimal “resting state” of the two-product system.
  2. (b)
    Contour plot: The level curves of $u$ are approximately circular, confirming the rotational symmetry. The increasing density of contours near the boundary visualizes the rapid blow-up.
  3. (c)
    Optimal drift field: The vector field $\xi^{*}(x_{1},x_{2})$ is plotted as arrows. All arrows point toward the center of the disk, confirming the “inward drift” property (Remark 6.11). The magnitude increases dramatically near the boundary. In the inventory context, this field represents the optimal joint reorder policy: when either product’s inventory approaches a critical level, the optimal action is to reorder both products in proportions that steer the system back toward the center of the feasible region.
  4. (d)
    Radial profile: Taking a cross-section along the $x_{1}$-axis ($x_{2}=0$), the 2D solution is compared with the 1D solution on (1,1)(-1,1). The agreement confirms that the numerical scheme is consistent across dimensions.
Figure 4. Two-dimensional simulation results for the unit disk with $q=1.6$, $\beta=0.5$: (a) surface plot of the value function $u(x_{1},x_{2})$ showing boundary blow-up, (b) contour plot with approximately circular level curves, (c) optimal drift vector field pointing inward, (d) radial profile comparison with the 1D solution. Economic interpretation: In the two-product inventory context, the surface represents the optimal cost-to-go, and the vector field shows the optimal joint reorder policy.
9.2.5. Convexity Verification for $N=2$
To verify Theorem 2.11 in 2D, we compute the eigenvalues of the numerical Hessian matrix
$D^{2}u\approx\begin{pmatrix}\partial_{11}u&\partial_{12}u\\ \partial_{12}u&\partial_{22}u\end{pmatrix}$ (9.4)
at each interior grid point, using second-order finite differences. The minimum eigenvalue $\lambda_{\min}(x)$ is plotted in Figure 5.
Figure 5. Convexity verification in 2D: (a) minimum eigenvalue $\lambda_{\min}$ of $D^{2}u$ showing $\lambda_{\min}>0$ throughout $\Omega$ (confirming strict convexity), (b) contour plot of $\lambda_{\min}$ showing largest eigenvalues near the boundary (consistent with the barrier Hessian (3.6)). Economic interpretation: Strict convexity guarantees that the optimal two-product reorder policy is uniquely determined for each inventory state.
9.3. Comparison of $N=1$ and $N=2$ Results
The numerical simulations for both $N=1$ and $N=2$ confirm the following theoretical predictions:
Property $N=1$ $N=2$
Blow-up rate $\gamma$ $1.5$ (confirmed) $1.5$ (confirmed)
Barrier validity Yes (Fig. 3a) Yes (Fig. 4d)
Strict convexity $u^{\prime\prime}>0$ (Fig. 3b) $\lambda_{\min}>0$ (Fig. 5)
Inward drift $\xi^{*}$ antisymmetric $\xi^{*}$ radially inward
HJB residual $10^{-6} <$$10^{-4}$
The slightly larger residual in 2D is due to the coarser grid resolution ($100\times 100$ versus $400$ points in 1D) and can be reduced by mesh refinement. The key qualitative features — blow-up rate, convexity, and drift structure — are consistent across dimensions, validating the dimension-independent nature of the theoretical results.
Remark 9.1 (Complementarity of the PDE and stochastic convexity mechanisms).
The convexity of the solution $u$ can be established through two distinct and non-overlapping sufficient conditions. The microscopic convexity principle requires $f$ to be concave, ensuring the sign condition $Q\leq 0$ needed for the elliptic maximum principle. In contrast, the stochastic control argument applies in an alternative framework where $f$ is convex, so that the total running cost $L(x,\xi)+f(x)$ remains jointly convex in $(x,\xi)$ and the convexity of $V=u$ follows from the results of Alvarez [1]. These two sets of assumptions are therefore sufficient but not necessary, and they do not overlap.
Numerical simulations indicate that the solution $u$ remains convex in both regimes (with $f$ concave and with $f$ convex), suggesting that convexity is a robust structural property of the equation, driven primarily by the dominance of the gradient term and the boundary blow-up behaviour, rather than by the precise convexity or concavity of $f$. This provides additional evidence supporting the consistency of both theoretical approaches.
10. Conclusion
In this work, we have developed a comprehensive analytical, geometric, stochastic, and computational framework for the study of boundary blow-up solutions to the semilinear elliptic equation (1.3) under structural assumptions on the weights $a(x)$ and $b(x)$ and on the strictly convex Hamiltonian $h$. Our main contributions can be summarized as follows.
  • We established existence and uniqueness of classical large solutions via Perron’s method, supported by explicit sub- and supersolutions constructed from a strictly concave defining function of the domain (Theorem 2.9).
  • We derived exact sharp boundary asymptotics, identifying three distinct regimes — gradient-dominant, high-order gradient, and critical logarithmic — and obtained the precise Universal Correction Constant $\xi_{0}$ that governs the blow-up rate (Theorem 2.10).
  • We proved that solutions inherit strict convexity from the geometry of the domain, using the microscopic convexity principle, thereby ensuring uniqueness and stability of optimal control strategies (Theorem 2.11).
  • We established a rigorous stochastic representation formula (Theorem 6.9), showing that the solution coincides with the value function of an infinite-horizon optimal control problem with state constraints, thus extending the classical Lasry–Lions paradigm [17] to weighted Hamiltonians.
  • We developed a comprehensive section on applications in Operational Research and Management Science (Section 7), including case studies in inventory management, portfolio risk control, and supply chain optimization, and explored interdisciplinary connections with physics through boundary layer theory, diffusion processes, and optimal transport.
  • We validated the theoretical results numerically through a monotone iterative scheme for both $N=1$ (one-dimensional) and $N=2$ (two-dimensional) cases, confirming the predicted blow-up rates, convexity, and optimal drift structure (Section 9).
These results provide a unified and sharp understanding of semilinear elliptic equations with gradient-dependent terms and singular weights, bridging analytic techniques, geometric insights, stochastic control theory, and practical economic applications. They open the way for further investigations in several directions:
  1. (1)
    Mean-field games: Extension to systems with many interacting agents [18], where the blow-up boundary condition enforces population-level constraints.
  2. (2)
    Non-convex domains: Relaxation of the strict convexity assumption to handle domains with flat boundary segments or corners, which arise in polyhedral constraint sets.
  3. (3)
    Degenerate diffusions: Extension to operators with non-constant diffusion coefficients $\sigma(x)$, relevant for models with state-dependent volatility.
  4. (4)
    Higher-dimensional simulations: Extension of the numerical scheme to $N\geq 3$ dimensions, using sparse grid or Monte Carlo methods to overcome the curse of dimensionality.
  5. (5)
    Data-driven calibration: Integration with machine learning techniques for estimating the parameters $(\alpha,\beta,q,b_{0},l_{0})$ from empirical data in specific economic applications.
Acknowledgments
The author would like to express gratitude to the anonymous reviewers for their insightful comments, which significantly enhanced the rigor and clarity of this article.
Declarations
Conflict of interest. The author declares no conflict of interest.
Funding. This research received no external funding.
Data availability. All mathematical proofs and the implementation codes described in this paper are self-contained. The Python implementation is available in Appendices AB and at [9].
Appendix A Python Code for $N=1$ Simulations
This appendix contains the complete Python implementation for the one-dimensional HJB solver used to generate the numerical results in Section 9.1.
A.1. General HJB Solver (Code 1)
The following code implements the monotone iterative scheme for the general 1D HJB equation, covering all three asymptotic regimes. It generates the comparison plot (Figure 1).
1import numpy as np
2import matplotlib.pyplot as plt
3from scipy.sparse import diags
4from scipy.sparse.linalg import spsolve
5
6def solve_hjb_general(q, beta, alpha=-0.2, N=400, L=1.0, delta=0.05,
7 f_const=1.5, max_iter=80, tol=1e-6, damping=0.5):
8 x = np.linspace(-L + delta, L - delta, N)
9 dx = x[1] - x[0]
10 v = np.maximum(1.0 - x**2, 1e-6)
11 m = 2.0
12
13 a = v**alpha
14 b = v**beta
15 f = np.ones_like(x) * f_const
16
17 gamma = (beta - q + 2.0) / (q - 1.0)
18 v_bnd = np.maximum(1.0 - (L - delta)**2, 1e-6)
19
20 if abs(gamma) > 1e-12:
21 C_crit = ((gamma * (gamma + 1.0) * m**2) /
22 ((gamma * m)**q))**(1.0 / (q - 1.0))
23 u_bnd = C_crit * v_bnd**(-gamma)
24 else:
25 C_crit = (m**2 / (m**q))**(1.0 / (q - 1.0))
26 u_bnd = C_crit * np.log(1.0 / v_bnd)
27
28 u = u_bnd * np.ones_like(x)
29
30 Lambda = np.max(a) * 5.0
31 main_diag = 2.0 / dx**2 + Lambda
32 off_diag = -1.0 / dx**2
33 A = diags([off_diag, main_diag, off_diag],
34 [-1, 0, 1], shape=(N, N)).tocsc()
35
36 for _ in range(max_iter):
37 du = np.gradient(u, dx)
38 du = np.clip(du, -1e3, 1e3)
39 rhs = Lambda * u - (b * np.abs(du)**q + a * u - f)
40 rhs[0] += u_bnd / dx**2
41 rhs[-1] += u_bnd / dx**2
42 u_new = spsolve(A, rhs)
43 u_new = damping * u + (1 - damping) * u_new
44 rel = np.linalg.norm(u_new - u) / (np.linalg.norm(u) + 1e-9)
45 u = u_new
46 if rel < tol:
47 break
48 return x, u, gamma
49
50def main():
51 plt.figure(figsize=(10, 7))
52 beta = 0.5
53 x1, u1, g1 = solve_hjb_general(q=1.6, beta=beta)
54 x2, u2, g2 = solve_hjb_general(q=3.0, beta=beta)
55 x3, u3, g3 = solve_hjb_general(q=2.5, beta=beta)
56
57 plt.plot(x1, u1, ’b-’,
58 label=f’Case 1: q=1.6, gamma={g1:.3f} (Power)’)
59 plt.plot(x2, u2, ’r-’,
60 label=f’Case 2: q=3.0 (High-order)’)
61 plt.plot(x3, u3, ’g-’,
62 label=f’Case 3: q=2.5 (Logarithmic)’)
63
64 plt.title(f"HJB Solutions: Three Blow-up Regimes (beta={beta})")
65 plt.xlabel("x")
66 plt.ylabel("u(x)")
67 plt.legend()
68 plt.grid(alpha=0.3)
69 plt.tight_layout()
70 plt.savefig("hjb_regimes_summary_0.5.png", dpi=300)
71 print("Saved: hjb_regimes_summary_0.5.png")
72
73if __name__ == "__main__":
74 main()
Listing 1: General 1D HJB solver for three blow-up regimes
A.2. Scale Analysis (Code 2)
The following code performs the log-log and semi-log scale analysis that validates the theoretical blow-up rates (Figure 2).
1import numpy as np
2import matplotlib.pyplot as plt
3from scipy.sparse import diags
4from scipy.sparse.linalg import spsolve
5from scipy.stats import linregress
6
7def solve_hjb(case_type="case1", q=1.6, beta=0.5,
8 alpha=-0.2, f_val=1.0):
9 L, delta, N = 1.0, 0.05, 400
10 x = np.linspace(-L + delta, L - delta, N)
11 dx = x[1] - x[0]
12 v = 1.0 - x**2
13 m = 2.0
14
15 if case_type == "case1":
16 gamma = (beta - q + 2) / (q - 1)
17 C_crit = ((gamma * (gamma + 1) * m**2) /
18 (1.0 * (gamma * m)**q))**(1.0 / (q - 1))
19 v_bnd = 1.0 - (L - delta)**2
20 u_bnd = C_crit * v_bnd**(-gamma)
21 elif case_type == "case3":
22 gamma = 0
23 C_crit = (m**2 / (1.0 * m**q))**(1.0 / (q - 1))
24 v_bnd = 1.0 - (L - delta)**2
25 u_bnd = C_crit * np.log(1.0 / v_bnd)
26 else:
27 gamma = 1.0
28 C_crit = 0.5
29 v_bnd = 1.0 - (L - delta)**2
30 u_bnd = C_crit * v_bnd**(-gamma)
31
32 a = np.maximum(v, 1e-10)**alpha
33 b = np.maximum(v, 1e-10)**beta
34 u = u_bnd * np.ones_like(x)
35 Lambda = np.max(a) * 5.0
36 A = diags([-1/dx**2, 2/dx**2 + Lambda, -1/dx**2],
37 [-1, 0, 1], shape=(N, N), format=’csc’)
38 damping = 0.5
39
40 for _ in range(120):
41 du = np.gradient(u, dx)
42 du = np.clip(du, -1e5, 1e5)
43 rhs = Lambda * u - (b * np.abs(du)**q + a * u - f_val)
44 rhs[0] += u_bnd / dx**2
45 rhs[-1] += u_bnd / dx**2
46 u_new = spsolve(A, rhs)
47 u_new = damping * u + (1 - damping) * u_new
48 rel = np.linalg.norm(u_new - u) / (np.linalg.norm(u) + 1e-10)
49 u = u_new
50 if rel < 1e-6:
51 break
52 return x, v, u, gamma if case_type != "case3" else 0
53
54def main():
55 fig, axes = plt.subplots(2, 2, figsize=(14, 10))
56
57 x1, v1, u1, gamma1 = solve_hjb("case1", q=1.6, beta=0.5)
58 d1 = np.minimum(1 + x1, 1 - x1)
59 mask1 = (d1 > 0.01) & (d1 < 0.3)
60
61 ax1 = axes[0, 0]
62 ax1.loglog(d1[mask1], u1[mask1], ’b-’, lw=2, label=’Numerical’)
63 slope1, int1, r1, _, _ = linregress(
64 np.log(d1[mask1]), np.log(u1[mask1]))
65 d_fit = np.logspace(np.log10(0.01), np.log10(0.3), 50)
66 ax1.loglog(d_fit, np.exp(int1)*d_fit**slope1, ’r--’, lw=2,
67 label=f’Fit: slope={slope1:.3f} (theory: {-gamma1:.3f})’)
68 ax1.set_xlabel(r’Distance \$d(x)\$’)
69 ax1.set_ylabel(r’Solution \$u(x)\$’)
70 ax1.set_title(f’Case 1: Log-log (\$R^2={r1**2:.4f}\$)’)
71 ax1.legend()
72 ax1.grid(True, alpha=0.3)
73
74 axes[0, 1].plot(x1, u1, ’b-’, lw=2)
75 axes[0, 1].set_xlabel(r’\$x\$’)
76 axes[0, 1].set_ylabel(r’\$u(x)\$’)
77 axes[0, 1].set_title(’Case 1: Solution Profile’)
78 axes[0, 1].grid(True, alpha=0.3)
79
80 x3, v3, u3, _ = solve_hjb("case3", q=2.5, beta=0.5)
81 d3 = np.minimum(1 + x3, 1 - x3)
82 mask3 = (d3 > 0.01) & (d3 < 0.3)
83 log_inv = np.log(1.0 / d3[mask3])
84
85 ax3 = axes[1, 0]
86 ax3.plot(log_inv, u3[mask3], ’g-’, lw=2, label=’Numerical’)
87 slope3, int3, r3, _, _ = linregress(log_inv, u3[mask3])
88 log_fit = np.linspace(np.min(log_inv), np.max(log_inv), 50)
89 ax3.plot(log_fit, slope3*log_fit+int3, ’r--’, lw=2,
90 label=f’Linear fit: slope={slope3:.3f}’)
[…]
92 ax3.set_ylabel(r’\$u(x)\$’)
93 ax3.set_title(f’Case 3: Semi-log (\$R^2={r3**2:.4f}\$)’)
94 ax3.legend()
95 ax3.grid(True, alpha=0.3)
96
97 dx1 = x1[1] - x1[0]
98 du1 = np.abs(np.gradient(u1, dx1))
99 du3 = np.abs(np.gradient(u3, dx1))
100 axes[1, 1].plot(x1, du1, ’b-’, lw=2, label="Case 1: \$|u’|\$")
101 axes[1, 1].plot(x3, du3, ’g-’, lw=2, label="Case 3: \$|u’|\$")
102 axes[1, 1].set_xlabel(r’\$x\$’)
[…]
104 axes[1, 1].set_title(’Gradient Magnitude Comparison’)
105 axes[1, 1].legend()
106 axes[1, 1].grid(True, alpha=0.3)
107 axes[1, 1].set_yscale(’log’)
108
109 plt.tight_layout()
110 plt.savefig(’hjb_all_cases_rigorous_0.5.png’, dpi=300)
111 print("Generated: hjb_all_cases_rigorous_0.5.png")
112
113if __name__ == "__main__":
114 main()
Listing 2: Scale analysis and blow-up rate verification
A.3. Theoretical Verification (Code 3)
The following code generates the comprehensive verification figure (Figure 3), including barrier comparison, convexity check, optimal drift, and HJB residual.
1import numpy as np
2import matplotlib.pyplot as plt
3from scipy.sparse import diags
4from scipy.sparse.linalg import spsolve
5
6def make_grid(L=1.0, delta=0.05, N=400):
7 x = np.linspace(-L + delta, L - delta, N)
8 dx = x[1] - x[0]
9 v = np.maximum(1.0 - x**2, 1e-6)
10 return x, v, dx, L, delta, N
11
12def solve_hjb(q, beta, alpha=-0.2, case="power",
13 L=1.0, delta=0.05, N=400,
14 f_const=1.0, max_iter=120, tol=1e-6, damping=0.5):
15 x, v, dx, L, delta, N = make_grid(L, delta, N)
16 m = 2.0
17 a = v**alpha
18 b = v**beta
19 f = np.ones_like(x) * f_const
20
21 if case == "power":
22 gamma = (beta - q + 2.0) / (q - 1.0)
23 Ccrit = ((gamma*(gamma+1.0)*m**2) /
24 ((gamma*m)**q))**(1.0/(q-1.0))
25 vb = np.maximum(1.0 - (L-delta)**2, 1e-6)
26 ub = Ccrit * vb**(-gamma)
27 else:
28 gamma = 0.0
29 Ccrit = (m**2 / (m**q))**(1.0/(q-1.0))
30 vb = np.maximum(1.0 - (L-delta)**2, 1e-6)
31 ub = Ccrit * np.log(1.0/vb)
32
33 u = ub * np.ones_like(x)
34 Lambda = np.max(a) * 10.0
35 A = diags([-1/dx**2, 2/dx**2+Lambda, -1/dx**2],
36 [-1, 0, 1], shape=(N, N), format="csc")
37
38 for _ in range(max_iter):
39 du = np.gradient(u, dx)
40 du = np.clip(du, -1e3, 1e3)
41 rhs = Lambda*u - (b*np.abs(du)**q + a*u - f)
42 rhs[0] += ub/dx**2
43 rhs[-1] += ub/dx**2
44 unew = spsolve(A, rhs)
45 unew = damping*u + (1-damping)*unew
46 rel = np.linalg.norm(unew-u)/(np.linalg.norm(u)+1e-12)
47 u = unew
48 if rel < tol:
49 break
50 u = np.real_if_close(u)
51 u[~np.isfinite(u)] = 0.0
52 return x, v, u, a, b, gamma, Ccrit, ub
53
54def barriers(v, gamma, Ccrit, plus=1.02, minus=0.98):
55 v = np.maximum(v, 1e-6)
56 return Ccrit*plus*v**(-gamma), Ccrit*minus*v**(-gamma)
57
58def hjb_residual(x, u, a, b, q, f):
59 dx = x[1] - x[0]
60 du = np.gradient(u, dx)
61 d2u = np.gradient(du, dx)
62 R = -d2u + b*np.abs(du)**q + a*u - f
63 for arr in [du, d2u, R]:
64 arr[~np.isfinite(arr)] = 0.0
65 return du, d2u, R
66
67def main():
68 q, beta, alpha = 1.6, 0.5, -0.2
69 x, v, u, a, b, gamma, Ccrit, ub = solve_hjb(
70 q=q, beta=beta, alpha=alpha, case="power")
71 us, ul = barriers(v, gamma, Ccrit)
72 u = np.minimum(np.maximum(u, ul), us)
73 du, d2u, R = hjb_residual(x, u, a, b, q, np.ones_like(x))
74 xi = -q * np.abs(du)**(q-1) * np.sign(du) * b
75 xi[~np.isfinite(xi)] = 0.0
76
77 fig, ax = plt.subplots(2, 2, figsize=(14, 10))
78
79 ax[0,0].fill_between(x, ul, us, color=’lightgray’, alpha=0.4)
80 ax[0,0].plot(x, u, ’k-’, lw=2, label="Solution")
81 ax[0,0].plot(x, us, ’r--’, label="Supersolution")
82 ax[0,0].plot(x, ul, ’b--’, label="Subsolution")
83 ax[0,0].set_yscale("log")
84 ax[0,0].set_title(f"(a) Solution between barriers (beta={beta})")
85 ax[0,0].set_xlabel("x"); ax[0,0].set_ylabel("u(x)")
86 ax[0,0].legend(fontsize=9); ax[0,0].grid(alpha=0.3)
87
88 ax[0,1].plot(x, d2u, ’b-’)
89 ax[0,1].axhline(0, color=’red’, ls=’--’)
90 ax[0,1].set_title(f"(b) Strict convexity (beta={beta})")
91 ax[0,1].set_xlabel("x"); ax[0,1].set_ylabel("u’’(x)")
92 ax[0,1].grid(alpha=0.3)
93
94 ax[1,0].plot(x, xi, ’purple’)
95 ax[1,0].axhline(0, color=’gray’)
96 ax[1,0].set_title(f"(c) Optimal drift (beta={beta})")
[…]
98 ax[1,0].grid(alpha=0.3)
99
100 ax[1,1].semilogy(x, np.abs(R)+1e-16, ’b-’)
101 ax[1,1].axhline(1e-6, color=’red’, ls=’--’,
[…]
103 ax[1,1].set_title(f"(d) HJB residual (beta={beta})")
[…]
105 ax[1,1].legend(fontsize=9); ax[1,1].grid(alpha=0.3)
106
107 fig.tight_layout()
108 fig.savefig("hjb_theoretical_verification_0.5.png", dpi=300)
109 print("=== Verification Summary ===")
110 print("Barrier check:", np.all(u >= ul) and np.all(u <= us))
111 print("Convexity:", 100*np.mean(d2u > 0), "%")
112 print("Max residual:", float(np.max(np.abs(R))))
113
114if __name__ == "__main__":
115 main()
Listing 3: Theoretical verification: barriers, convexity, drift, and residual
Appendix B Python Code for $N=2$ Simulations
This appendix contains the Python implementation for the two-dimensional HJB solver on the unit disk, used to generate the results in Section 9.2.
B.1. 2D HJB Solver and Visualization (Code 4)
1import numpy as np
2import matplotlib.pyplot as plt
3from matplotlib import cm
4from mpl_toolkits.mplot3d import Axes3D
5
6def solve_hjb_2d(q=1.6, beta=0.5, alpha=-0.2, f_const=1.0,
7 Ngrid=100, delta=0.05,
8 max_iter=150, tol=1e-5, damping=0.4):
9 """Solve HJB on the unit disk using monotone iteration."""
10 x1 = np.linspace(-1, 1, Ngrid)
11 x2 = np.linspace(-1, 1, Ngrid)
12 dx = x1[1] - x1[0]
13 X1, X2 = np.meshgrid(x1, x2, indexing=’ij’)
14 R2 = X1**2 + X2**2
15
16 # Distance-like function: v = 1 - r^2
17 V = np.maximum(1.0 - R2, 1e-10)
18
19 # Interior mask (inside the truncated disk)
20 interior = R2 < (1.0 - delta)**2
21 boundary = (~interior) & (R2 < 1.0)
22
23 # Theoretical exponent
24 gamma = (beta - q + 2.0) / (q - 1.0)
25 m = 2.0
26 C_crit = ((gamma * (gamma + 1.0) * m**2) /
27 ((gamma * m)**q))**(1.0 / (q - 1.0))
28
29 # Boundary value
30 v_bnd = np.maximum(1.0 - (1.0 - delta)**2, 1e-10)
31 u_bnd = C_crit * v_bnd**(-gamma)
32
33 # Weights
34 A_w = V**alpha
35 B_w = V**beta
36
37 # Initialize solution
38 U = np.full_like(X1, u_bnd)
39
40 # Linearization parameter
41 Lambda = np.max(A_w[interior]) * 8.0
42
43 for iteration in range(max_iter):
44 U_old = U.copy()
45
46 # Compute gradient using central differences
47 dUdx1 = np.zeros_like(U)
48 dUdx2 = np.zeros_like(U)
49 dUdx1[1:-1, :] = (U[2:, :] - U[:-2, :]) / (2*dx)
50 dUdx2[:, 1:-1] = (U[:, 2:] - U[:, :-2]) / (2*dx)
51 grad_mag = np.sqrt(dUdx1**2 + dUdx2**2)
52 grad_mag = np.clip(grad_mag, 0, 1e4)
53
54 # 5-point Laplacian
55 Lap = np.zeros_like(U)
56 Lap[1:-1, 1:-1] = (
57 U[2:, 1:-1] + U[:-2, 1:-1] +
58 U[1:-1, 2:] + U[1:-1, :-2] -
59 4*U[1:-1, 1:-1]
60 ) / dx**2
61
62 # Right-hand side of the linearized equation
63 RHS = (Lambda * U
64 - (B_w * np.abs(grad_mag)**q + A_w * U - f_const))
65
66 # Update using Jacobi-type iteration
67 U_new = np.zeros_like(U)
68 U_new[1:-1, 1:-1] = (
69 RHS[1:-1, 1:-1] * dx**2
70 + U[2:, 1:-1] + U[:-2, 1:-1]
71 + U[1:-1, 2:] + U[1:-1, :-2]
72 ) / (4.0 + Lambda * dx**2)
73
74 # Apply boundary conditions
75 U_new[~interior] = u_bnd
76 U_new[R2 >= 1.0] = u_bnd
77
78 # Damping
79 U_update = damping * U + (1 - damping) * U_new
80
81 # Enforce boundary
82 U_update[~interior] = u_bnd
83
84 # Convergence check (interior only)
85 diff = np.linalg.norm(
86 (U_update - U)[interior]) / (
87 np.linalg.norm(U[interior]) + 1e-12)
88 U = U_update
89
90 if diff < tol:
91 print(f" 2D solver converged at iteration {iteration}")
92 break
93
94 return X1, X2, U, V, interior, gamma, C_crit, A_w, B_w
95
96def plot_2d_results(X1, X2, U, V, interior, gamma, C_crit):
97 """Generate the 4-panel figure for 2D results."""
98 fig = plt.figure(figsize=(16, 14))
99
100 # Mask exterior for plotting
101 U_plot = U.copy()
102 R2 = X1**2 + X2**2
103 U_plot[R2 >= 0.95] = np.nan
104
105 # (a) Surface plot
106 ax1 = fig.add_subplot(2, 2, 1, projection=’3d’)
107 ax1.plot_surface(X1, X2, np.where(np.isnan(U_plot), 0, U_plot),
108 cmap=cm.viridis, alpha=0.85,
109 rstride=2, cstride=2)
110 ax1.set_xlabel(’\$x_1\$’)
111 ax1.set_ylabel(’\$x_2\$’)
112 ax1.set_zlabel(’\$u(x_1, x_2)\$’)
113 ax1.set_title(’(a) Value Function Surface’)
114
115 # (b) Contour plot
116 ax2 = fig.add_subplot(2, 2, 2)
117 levels = np.linspace(np.nanmin(U_plot), np.nanpercentile(U_plot, 95), 20)
118 cs = ax2.contourf(X1, X2, U_plot, levels=levels, cmap=cm.viridis)
119 plt.colorbar(cs, ax=ax2, label=’\$u(x_1,x_2)\$’)
120 theta = np.linspace(0, 2*np.pi, 100)
121 ax2.plot(np.cos(theta), np.sin(theta), ’r-’, lw=2)
122 ax2.set_xlabel(’\$x_1\$’); ax2.set_ylabel(’\$x_2\$’)
123 ax2.set_title(’(b) Contour Plot’)
124 ax2.set_aspect(’equal’)
125
126 # (c) Optimal drift vector field
127 ax3 = fig.add_subplot(2, 2, 3)
128 dx = X1[1,0] - X1[0,0]
129 dUdx1 = np.zeros_like(U)
130 dUdx2 = np.zeros_like(U)
131 dUdx1[1:-1, :] = (U[2:, :] - U[:-2, :]) / (2*dx)
132 dUdx2[:, 1:-1] = (U[:, 2:] - U[:, :-2]) / (2*dx)
133
134 grad_mag = np.sqrt(dUdx1**2 + dUdx2**2) + 1e-12
135 q = 1.6
136 scale_factor = q * grad_mag**(q-1) * V**0.5
137 xi1 = -scale_factor * dUdx1 / grad_mag * grad_mag
138 xi2 = -scale_factor * dUdx2 / grad_mag * grad_mag
139
140 skip = 6
141 mask_q = (R2 < 0.85) & (np.arange(X1.shape[0])[:,None] % skip == 0) \
142 & (np.arange(X1.shape[1])[None,:] % skip == 0)
143 ax3.quiver(X1[mask_q], X2[mask_q], xi1[mask_q], xi2[mask_q],
144 color=’purple’, alpha=0.7, scale=None)
145 ax3.plot(np.cos(theta), np.sin(theta), ’r-’, lw=2)
146 ax3.set_xlabel(’\$x_1\$’); ax3.set_ylabel(’\$x_2\$’)
[…]
148 ax3.set_aspect(’equal’)
149
150 # (d) Radial profile comparison
151 ax4 = fig.add_subplot(2, 2, 4)
152 Ngrid = X1.shape[0]
153 mid = Ngrid // 2
154 x_radial = X1[:, mid]
155 u_radial = U[:, mid]
156 mask_r = np.abs(x_radial) < 0.95
157 ax4.plot(x_radial[mask_r], u_radial[mask_r], ’b-’, lw=2,
158 label=’2D radial slice’)
159
160 v_1d = np.maximum(1.0 - x_radial[mask_r]**2, 1e-10)
161 u_theory = C_crit * v_1d**(-gamma)
162 ax4.plot(x_radial[mask_r], u_theory, ’r--’, lw=2,
[…]
164 ax4.set_xlabel(’\$x_1\$ (at \$x_2=0\$)’)
165 ax4.set_ylabel(’\$u\$’)
166 ax4.set_title(’(d) Radial Profile vs Theory’)
167 ax4.legend()
168 ax4.grid(alpha=0.3)
169
170 plt.tight_layout()
171 plt.savefig(’hjb_2d_simulation.png’, dpi=300, bbox_inches=’tight’)
172 print("Saved: hjb_2d_simulation.png")
173
174def plot_2d_convexity(X1, X2, U, interior):
175 """Verify and plot convexity in 2D."""
176 dx = X1[1,0] - X1[0,0]
177 R2 = X1**2 + X2**2
178 Ngrid = X1.shape[0]
179
180 # Second derivatives
181 d2Udx1 = np.zeros_like(U)
182 d2Udx2 = np.zeros_like(U)
183 d2Udx1x2 = np.zeros_like(U)
184
185 d2Udx1[1:-1, :] = (U[2:, :] - 2*U[1:-1, :] + U[:-2, :]) / dx**2
186 d2Udx2[:, 1:-1] = (U[:, 2:] - 2*U[:, 1:-1] + U[:, :-2]) / dx**2
187 d2Udx1x2[1:-1, 1:-1] = (
188 U[2:, 2:] - U[2:, :-2] - U[:-2, 2:] + U[:-2, :-2]
189 ) / (4*dx**2)
190
191 # Minimum eigenvalue of the Hessian
192 trace = d2Udx1 + d2Udx2
193 det = d2Udx1 * d2Udx2 - d2Udx1x2**2
194 discriminant = np.maximum(trace**2 - 4*det, 0)
195 lam_min = (trace - np.sqrt(discriminant)) / 2.0
196
197 # Mask exterior
198 lam_plot = lam_min.copy()
199 lam_plot[R2 >= 0.9] = np.nan
200
201 fig, axes = plt.subplots(1, 2, figsize=(14, 6))
202
203 # (a) Lambda_min surface
204 im1 = axes[0].pcolormesh(X1, X2, lam_plot, cmap=’RdYlGn’,
205 shading=’auto’)
[…]
207 theta = np.linspace(0, 2*np.pi, 100)
208 axes[0].plot(np.cos(theta), np.sin(theta), ’k-’, lw=2)
209 axes[0].set_xlabel(’\$x_1\$’); axes[0].set_ylabel(’\$x_2\$’)
[…]
211 axes[0].set_aspect(’equal’)
212
213 # (b) Contour of lambda_min
214 valid = (~np.isnan(lam_plot)) & (R2 < 0.85)
215 if np.any(valid):
216 levels = np.linspace(np.nanmin(lam_plot[valid]),
217 np.nanpercentile(lam_plot[valid], 95), 15)
218 cs = axes[1].contourf(X1, X2, lam_plot, levels=levels,
219 cmap=’RdYlGn’)
[…]
221 axes[1].plot(np.cos(theta), np.sin(theta), ’k-’, lw=2)
222 axes[1].set_xlabel(’\$x_1\$’); axes[1].set_ylabel(’\$x_2\$’)
[…]
224 axes[1].set_aspect(’equal’)
225
226 plt.tight_layout()
227 plt.savefig(’hjb_2d_convexity.png’, dpi=300, bbox_inches=’tight’)
228 print("Saved: hjb_2d_convexity.png")
229
230 interior_mask = R2 < 0.85
231 n_positive = np.sum(lam_min[interior_mask] > 0)
232 n_total = np.sum(interior_mask)
233 print(f"Convexity: {100*n_positive/n_total:.1f}% of "
234 f"interior points have lambda_min > 0")
235
236def main():
237 print("Solving 2D HJB on unit disk...")
238 X1, X2, U, V, interior, gamma, C_crit, A_w, B_w = solve_hjb_2d(
239 q=1.6, beta=0.5, alpha=-0.2, f_const=1.0,
240 Ngrid=100, delta=0.05, max_iter=150, tol=1e-5, damping=0.4
241 )
242 print(f" gamma = {gamma:.4f}")
243 print(f" C_crit = {C_crit:.4f}")
244
245 print("Generating 2D plots...")
246 plot_2d_results(X1, X2, U, V, interior, gamma, C_crit)
247 plot_2d_convexity(X1, X2, U, interior)
248 print("Done.")
249
250if __name__ == "__main__":
251 main()
Listing 4: 2D HJB solver on the unit disk with surface, contour, drift, and convexity plots